### v1

# Towards Aligning Multimodal LLMs with Human Experts: A Focus on Parent–Child Interaction

Abstract.
While multimodal large language models (MLLMs) are increasingly applied in human-centred AI systems, their ability to understand complex social interactions remains uncertain. We present an exploratory study on aligning MLLMs with speech–language pathologists (SLPs) in analysing joint attention in parent–child interactions, a key construct in early social–communicative development. Drawing on interviews and video annotations with three SLPs, we characterise how observational cues of gaze, action, and vocalisation inform their reasoning processes. We then test whether an MLLM can approximate this workflow through a two-stage prompting, separating observation from judgment. Our findings reveal that alignment is more robust at the observation layer, where experts share common descriptors, than at the judgement layer, where interpretive criteria diverge. We position this work as a case-based probe into expert–AI alignment in complex social behaviour, highlighting both the feasibility and the challenges of applying MLLMs to socially situated interaction analysis.
Parent-child interaction, Multimodal Large language model,Human-AI alignment, Speech language pathologist, Joint attention
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in collaborative and social computing
Figure 1. Conceptual overview: We investigate how MLLMs can be guided to align with SLPs in parent–child interaction through two key stages—Observation, using structured prompts based on behavioural cues such as gaze, action, and vocalisation; and Judgement, informed by few-shot learning with expert-derived practices.:
1. Introduction
Early parent-child interactions lay the foundation for lifelong communication, social, and cognitive development. Foundational social-communicative behaviours–such as joint attention, joint intention, and social referencing–emerge during play and serve as critical indicators of developmental progress. Yet these behaviours are often subtle and difficult for caregivers to identify without professional support. This challenge is consequential: approximately one in six children in the United States experiences at least one developmental delay (robinson2017cdc), with language delays among the most prevalent (sunderajan_speech_2019). At the same time, access to speech-language pathologists (SLPs)–professionals who assess children’s communicative development by attending to multimodal cues such as gaze, gesture, vocalisation, and context–remains limited, particularly in rural or underserved communities (o2005barriers). As a result, expert developmental knowledge and care often remain out of reach in everyday caregiving contexts.
Despite recent advances in artificial intelligence (AI) for family and education (yuan2024designing; sun2024exploring; fiani2024exploring; nikkhah2024family; currin2024opportunities; su2024hidden), existing systems primarily assist with creating therapy materials or scaffolding caregiver engagement, rather than emulating the observing and judging practices of SLPs themselves. For example, Lewis et al. (lewis2025exploring) investigated how SLPs interact with AI-generated materials for children from culturally and linguistically diverse backgrounds, and Dangol et al.(dangol2025want) investigated how AI might support parents in delivering therapy at home, highlighting emotional and logistical challenges. Yet in both cases, the evaluative expertise of SLPs remained external to the AI systems. Little is known about whether MLLMs can be aligned to interpret parent-child interactions in ways that reflect expert reasoning.
In this paper, we explore how multimodal large language models (MLLMs) can be guided to emulate SLPs’ observing and judging practices. We focus on joint attention as a representative case: a core developmental milestone that SLPs typically evaluate through holistic, experience-based interpretations of children’s gaze, actions, and vocalisations. While MLLMs excel at structured tasks such as video captioning or temporal grounding, they struggle with socially nuanced behaviours that lack clear ground truth. We therefore investigate how expert criteria can be translated into structured representations that allow MLLMs to more closely align with expert judgment.
Guided by this objective, we examine the following research questions (RQs):
  • RQ1: How do SLPs observe and analyse parent-child interactions? What criteria do they use to identify segments of strong or poor joint attention? How consistent are their judgments?
  • RQ2: How can we represent SLPs’ evaluative criteria in formats that MLLMs can both understand and execute, and what design choices best support alignment between MLLM and expert observation and judgement?
To this end, we conducted in-depth interviews and annotation sessions with three experienced SLPs, who analysed 25 videos of parent-child interactions for strong and poor instances of joint attention. From their annotations, we derived three key behavioural dimensions–gaze, action, and vocalisation–that underpin expert assessments. We then translated these heuristics into structured, MLLM-compatible task formats: first prompting models to extract behavioural segments from raw video, and then prompting them to assess interaction quality using few-shot examples.
Our key contributions are:
  • We contribute an account of expert judgement practices around joint attention in parent-child interactions. Drawing on interviews and annotation analyses with SLPs, we show that their evaluations hinge on three interrelated cues: gaze, action, and vocalisation.
  • We design and evaluate an exploratory MLLM system that aligns with speech-language pathologists’ approaches to joint attention assessment in two stages: (1) observing fine-grained behavioural cues from parent-child interaction videos using expert-informed prompting, achieving up to 85% accuracy across dimensions; and (2) evaluating interaction quality using only structured behavioural descriptions, reaching over 57% average accuracy compared to expert labels.
  • We curate a segment-level dataset with expert-labelled joint attention behaviours and derive a set of practical design guidelines for building MLLM-based systems that align SLP’s observing and judging process. These guidelines cover prompt engineering, cue structuring, model configuration, and future directions for parent-facing AI systems.
2. Related Work
2.1. Language Delay and the Limits of Current Parent Support
Language delay is one of the most common developmental concerns in early childhood, with prevalence estimates ranging from 2.3% to 19% among children aged 2 to 7 years (mclaughlin2011speech). Without timely intervention, children with language delay are at increased risk for later difficulties in reading, attention regulation, and social interaction (sunderajan2019speech). SLPs play a key role in identifying and supporting children with language delays through structured interventions. However, limited access to professional services–whether due to long wait times, high cost, or service shortages–has led to growing interest in empowering parents to support their children at home (lieneman2017parent; straitstimes2024earlyintervention).
Parent training programmes such as It Takes Two to Talk, More Than Words, and DIR Floortime have shown success in encouraging behaviours that promote early communication and social development (pepper2004talk; sussman1999words; brock2014statewide; dir_floortime). These interventions typically guide caregivers to be more responsive and attuned during everyday interactions, and have demonstrated benefits in improving the quantity and quality of parent-child communication. Yet while such programmes can teach parents strategies to encourage communication, they rarely equip caregivers to evaluate the developmental effectiveness of those interactions (finke2009all; gadberry2011survey; ganz2013impacts).
For many parents, identifying nuanced developmental milestones can be difficult without professional guidance. This is especially true for joint attention (joint attention), one of the earliest and most important indicators of social-communicative development (mundy2007individual; tomasello1986joint). Determining whether joint attention is strong, weak, or absent often requires the kind of interpretive expertise that SLPs develop over years of practice and is challenging to translate into formal rules (o2005barriers). Families outside of clinical environments rarely benefit from such expert judgment. Motivated by this gap, we undertook this exploratory study oh how MLLMs might be aligned with SLPs’ reasoning to assess joint attention in naturalistic parent-child interactions. By probing the possibilities and limitations of this approach, we seek to inform future efforts to design AI tools that can augment professional expertise and broaden access to developmental support.
2.2. Technological Support for Parent-Child Interaction
There has been a surge in interactive technologies designed to support early childhood development, particularly by enhancing the quality and frequency of parent-child interaction. Early systems focused on lightweight support through manual logging. For instance, Chan et al. (chan2017wakey) developed WAKEY, which aimed to smooth morning routines by prompting parents to record schedules and track phrase usage. While effective in structuring communication, WAKEY relied entirely on manual input, without the ability to capture richer contextual signals.
Subsequent work shifted towards real-time speech analysis to provide parents with more immediate feedback. TalkBetter (hwang2014talkbetter) analysed turn-taking and generated tailored recommendations to scaffold language development, while TalkLIME (song2016talklime) extended this approach by offering live metrics such as utterance count, initiation ratio, and turn-taking frequency.
More recent systems incorporated multimodal sensing to deepen their understanding of interaction dynamics. Captivate! (kwon2022captivate) leveraged gaze estimation and speech recognition to detect joint attention and suggest parent prompts during play, and AACessTalk (choi2025aacess) supported minimally verbal autistic children by combining vocabulary recommendations with contextual guidance, ultimately improving turn-taking and parental reflection over a two-week deployment.
Parallel to these system developments, researchers have also explored how parents can be supported in adopting therapy strategies more effectively. Dangol et al. (dangol2025want) investigated how AI might address emotional and logistical barriers to at-home therapy, while Li et al. (li2025asd) introduced ASD-HI, a multimodal dataset and baseline model designed to detect strategies and assess fidelity in parent-led interventions.
Together, this trajectory reflects a shift from manual, parent-driven tools to increasingly automated, multimodal, and expert-aligned systems, yet few have investigated whether AI can evaluate interaction quality against professional standards–a gap our work explores by aligning MLLM-based analysis with SLP judgement.
2.3. Technological Support for SLPs
Generative AI is beginning to influence many aspects of clinical practice, including how SLPs source or adapt therapy materials. For example, Lewis et al. (lewis2025exploring) examined how SLPs interact with AI-generated visual content when working with children from culturally and linguistically diverse backgrounds. While their study highlighted important concerns regarding representation, bias, and contextual mismatch, it focused primarily on content usability. Their work does not address how AI might support or simulate the evaluative reasoning processes that SLPs use when observing and interpreting children’s behaviours in real time. With the rapid progress of MLLMs across tasks such as video captioning and temporal grounding (lu2024gpt), researchers have begun to explore their potential for analysing human behaviour–particularly in socially grounded contexts such as parent-child interaction. MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (gandhi2023multimodal; jain2024vcoder). Thanks to their in-context learning capabilities, MLLMs can flexibly adapt to new interaction settings with minimal data (dos2023composite; whitehead2024generative), making them promising for contexts that require domain-specific reasoning. Crucially, these models shift the output from fixed metrics or visualisation into natural language descriptions–potentially bridging the gap between raw behavioural signals and expert interpretation. Zheng et al.(zheng2024soap) introduced SOAP.AI, a system that enables experts to interact with MLLMs through in-context prompting and collaborative task design, and supports the automatic generation of initial behavioural segments–such as those relevant to joint attention–in domains like healthcare and education. However, SOAP.AI was designed primarily as a tool for expert users, without evaluating the reliability of its outputs or aligning them with SLP judgement. Extending this line of inquiry, Zheng et al.(zheng2025ai) subsequently explored how AI-generated documentation could more directly support SLPs, identifying four opportunities and proposing three fluidity-focused design goals for future systems. However, while SOAP.AI exemplifies how AI can augment expert workflows, it is designed primarily for expert users and does not evaluate the reliability of its outputs or align them with SLP judgement, leaving open the question of whether MLLMs can simulate expert evaluative capabilities.
While these approaches centre on supporting SLPs in their clinical workflows, they do not address a critical question: can current technologies simulate the observational and interpretive reasoning that SLPs apply when evaluating parent-child interactions? Bridging this gap requires rethinking how AI systems can align with expert judgement–not just to assist SLPs, but to empower parents with meaningful, expert-level insights into their everyday interactions. Our work explores the foundational alignment of MLLMs with SLP reasoning, laying the basis for more reliable support to both experts and caregivers.
2.4. Towards SLP-Aligned Observation and Judgement in Parent-Child Interaction
The first attempt to align MLLMs with the evaluative judgement of SLPs was introduced by Shi et al. (shi2025towards), who prompted MLLMs with definitions of joint attention to directly identify strong or poor joint attention segments through a temporal grounding task. Their results showed poor performance, primarily due to the models’ limited sensitivity to fine-grained gaze cues–particularly eye contact, which is central to joint attention dynamics. These findings highlight the need for closer alignment between MLLM interpretation and the evaluative reasoning used by SLPs, rather than relying solely on definition-based prompting.
Our work continues this exploration by examining how MLLMs can be guided toward more SLP-aligned interpretations of parent-child interaction.
3. Exploratory Study of SLPs’ Understanding of Parent-Child Interaction Based on Joint Attention
We began with a formative study to ground our exploration of how MLLMs might align with SLPs’ practices. First, we curated a corpus of 25 publicly available parent-child interaction videos from YouTube, which provided a diverse set of scenarios beyond the constraints of lab-based data collection. We then conducted semi-structured interviews with three experienced SLPs to probe how they conceptualise joint attention within Parent-Child Interaction and to surface the behavioural cues they attend to when judging interaction quality. Building on these discussions, the SLPs annotated selected video segments that exemplified both strong and poor instances of joint attention, producing rich, expert-grounded labels that served as training and evaluation data for our MLLM-based system. This study was reviewed and approved by our institution’s ethics board [number redacted for review], and all participants provided informed consent. SLPs received approximately USD 31.2 as compensation for participation.
3.1. Dataset
We collected videos from YouTube using targeted search terms (e.g., “parent-child interaction”, “play with child”, “play with baby”, “floor time with child”, and “talk to your baby”) to capture demonstration-based home-setting parent-child interaction sessions. Our goal was to find clear dyadic parent-child exchanges with continuous footage that flowed smoothly from beginning to end, allowing observation of how joint attention behaviours emerge and develop. Videos were required to provide clear behaviours, with stable framing and adequate audio, though minor background noise was acceptable. We excluded short or heavily edited clips and prioritised diverse, meaningful activities such as language learning, skill-building, and daily routines across a broad range of child ages. Elements such as subtitles or transitional animations were removed to preserve the raw parent-child interaction content. Figure 2 shows our video screening process. For the purposes of our study–aligning MLLM with human expertise, we noted that many of the selected videos were created by SLPs or professional organisations. Accordingly, in the subsequent stage, we recruited SLPs as our human experts.
Figure 2. Flow chart of the video screening process :
We curated a final dataset of 25 parent-child interaction videos (Video 1-25) spanning developmental stages from infancy to early school age (0-8 years), with most focusing on preschool years. Age information was obtained from video descriptions, covering children aged 0.5-2 (n=3), 2-4 (n=5), 4-6 (n=16), and 6-8 (n=1). Videos ranged from 30 seconds to 12 minutes, though most were short (13 under 1 minute, 11 between 1-5 minutes). A full list of video IDs, titles, and source URLs is provided in Appendix A. These videos were created and shared by parent-child interaction experts who have published numerous resources aimed at modelling effective interaction techniques and communication strategies for caregivers. Given the difficulty of accessing authentic recordings online and the exploratory nature of this study, most of our dataset consists of practitioner-shared demonstration sessions, often created by certified SLP or PCIT trainers. These videos were openly accessible and chosen with the intention of enhancing accessibility at this early stage.
The videos in the dataset covers three categories: (1) behavioural guidance and skill modelling [^1] https://www.youtube.com/watch?v=YUkujhg6j6w (n=10), where parents demonstrate techniques such as PRIDE skills (masse2018taking) or “Big Ignore” (woodfield2021time); (2) language and cognitive development [^2] https://www.youtube.com/watch?v=rVqJacvywAQ (n=9), showing tasks like topic discussions or Piaget’s experiments; and (3) daily life skills and interaction [^3] https://www.youtube.com/watch?v=N3wAPLXd7I0 (n=6), including reward chart reviews or shared playtime.
Examples from each category are illustrated in Figure 3.
(a) Behavioural Guidance and Skill Modelling: Parent uses the ”Big Ignore” (woodfield2021time) technique while the child continues painting (Video 3 (lieneman2023)).
(b) Language and Cognitive Development: Parent and child play a toy-hiding game to foster engagement. (Video1 10 (Lieneman2024))
(c) Daily Life Skills and Interaction: Parent and child build a marble run together, encouraging planning (Video2 (adam2013)).
Figure 3. Examples from three categories in our dataset: behavioural guidance, language development, and daily life interaction.:
3.2. Participants
Our three expert SLPs were all female, aged between 30 and 59 years (see Table 1). All have extensive clinical experience working directly with children, particularly those on the autism spectrum, and have received training in a wide range of evidence-based intervention programmes, including Hanen’s It Takes Two to Talk (pepper2004talk), More Than Words and DIR Floortime (dir_floortime).
Their clinical practice spans public and private sectors and includes diverse settings such as early intervention centres, preschools, home-based care, and multidisciplinary teams. Notably, the SLPs have worked in different countries across multiple continents, contributing to a culturally informed understanding of parent-child interaction.
Given their specialised training and long-standing experience, particularly with children on the autism spectrum, who often experience joint attention difficulties–the SLPs were well-qualified to identify and evaluate strong and poor instances of joint attention.
Table 1. Profiles of the three expert SLPs who participated in this study, including their age group, academic qualifications, specialist certifications, and experience years.
ID Age Group Qualifications Experience and Training
SLP1 30-34 BSc in Speech Pathology; Certified in Hanen (It Takes Two to Talk, More Than Words); DIR Floortime 7 years of paediatric therapy experience in early childhood communication support, primarily in homes and SLP centres
SLP2 30-34 BSc in Speech Pathology; Certified in Hanen (It Takes Two to Talk, More Than Words); DIR Floortime; PECS; PROMPT; Social Thinking; TalkTools; Key Word Sign Australia 9 years of paediatric therapy experience, including work with sensory processing and oral placement therapy approaches, primarily in SLP centres
SLP3 55-59 BSc in Speech Pathology; Msc in Communication Disorders; Certified in Hanen (It Takes Two to Talk, More Than Words); DIR Floortime 23 years of paediatric therapy experience, primarily in school settings, with extensive international practice experience
3.3. Study Protocol
We conducted our study with each SLP individually. The entire study was audio recorded. The study comprises three stages and took approximately 2 hours:
Pre-interview (10 minutes)
Each SLP was first asked to briefly share their professional background. They were then asked to describe what joint attention means in their professional judgment, what constituted strong, moderate, or poor joint attention, the behavioural cues that they typically attend to, and how contextual factors may influence their evaluations. They were also asked to share its significance for child development, and explain how joint attention typically emerges and is supported through Parent-Child Interaction.
Annotation of joint attention (80 minutes)
Each SLP then reviewed all 25 videos. To focus our analysis and ensure task feasibility, we limited annotations to the child’s joint attention behaviour, which is often a primary concern in early developmental assessment. They marked the start and end times of segments they considered to show strong or poor joint attention, with all unlabelled segments treated as moderate by default. Each annotation included their short explanation justifying their decision and their suggestions for how Parent-Child Interaction could be supported. We used our custom-built video annotation tool to facilitate the SLPs’ judgement process (see Figure 4, shown here with a screenshot from Video 24444https://www.youtube.com/watch?v=rVqJacvywAQ (adam2013)).
Post-task interview (20 minutes)
Finally, we revisited their annotation decisions to probe which observational cues (gaze, action, vocalisation) they relied on most, how they differentiated strong from poor instances.
Figure 4. Our video annotation tool supports SLPs’ judgement process with four main components: Video Playback Area for watching, pausing, and replaying parent–child interactions (shown here with a screenshot from Video 24. Timeline Annotation Area for selecting and labelling segments as strong or poor joint attention; Note-Taking Area for recording justifications or observations; and Control Button Area for task submission and navigation. :
3.4. Data Analysis
The audio recordings were transcribed using WhisperX (bain2023whisperx) and manually verified. We analysed the qualitative data by combining the SLPs’ transcripts from the interviews with their annotations during the video tasks. To facilitate comparison across experts and to quantify areas of agreement and divergence, we standardised the data by dividing each video into fixed 5-second segments and mapping SLP labels onto these units. Any segments not explicitly labelled as Strong or Poor were automatically assigned a Moderate label.
3.5. Interview Study Results: SLPs’ Conceptual Understanding of Joint Attention
In our semi-structured interviews, we sought to understand how each SLP conceptualises joint attention. We asked them to define what joint attention means in their clinical judgment, how they differentiate between strong and poor instances, and whether they expect developmental differences in how joint attention is expressed across child age groups. We summarised key insights from the interviews by grouping similar responses based on how the SLPs described and distinguished different forms of joint attention (see Table 2).
Table 2. SLPs’ conceptual definitions of joint attention, their evaluative criteria, and their expectations during play across developmental stages.
ID How They Define the Child’s Joint Attention What Counts as Strong vs. Poor How Developmental Stage Affects Expression How Joint Attention Appears During Play
SLP1 Joint attention is the child’s ability to share focus and intent with another person, not just look at the same thing. It’s about social connection. Strong: back-and-forth referencing, shared interest
Poor: no acknowledgment of other person’s presence
Yes – younger children often rely more on gesture, posture, or affect due to limited language and attention control. For older children, SLP1 expects clearer behavioural signals and social reciprocity. Joint attention fluctuates rapidly during play, especially in younger children. Distraction is expected and not inherently problematic. What matters is the overall balance–occasional poor moments are acceptable as long as strong episodes also occur.
SLP2 Joint attention is when a child follows or initiates shared gaze toward an object or action, and shows awareness that the other person is also attending. Strong: gaze shifts, alternating eye contact
Poor: no gaze response, ignores joint bids
Yes – early on, gaze behaviour may be less consistent. By age 3, SLP2 expects reliable referential gaze, as cognitive and visual coordination typically improve with development. Joint attention during play is highly variable and can shift within seconds. It’s normal for both strong and poor moments to co-occur within a short time window, such as within the same minute.
SLP3 Joint attention is a shared emotional and cognitive moment, often involving mutual engagement and affective alignment. It’s not just about behaviour but about intention. Strong: emotional reciprocity, coordinated interaction
Poor: child is disengaged despite cues
Yes – SLP3 accounts for language, motor, and emotional maturity. For younger children, subtle affective signals may be enough. For preschoolers, SLP3 looks for intentional behaviours like pointing or verbal expression. SLP3 believes it is unrealistic for a child to maintain strong joint attention throughout an entire play session. A child who always appears highly attentive–for example, staring continuously at the parent–would seem unusual. SLP3 described joint attention during play as something that “rises and falls like a temperature bar”, with natural shifts between strong and poor moments. What matters is the overall balance across the interaction, not perfection.
These interviews revealed that while all three SLPs shared a common understanding of joint attention as a socially coordinated process, they differed in how they conceptualised and prioritised specific aspects of it. SLP1 emphasised the coordination of multiple communicative behaviours, and vocalisation–within dynamic, reciprocal interactions between child and parent. SLP2 focused more narrowly on whether the child performed clear gaze shifts between the adult and the shared object, treating visual referencing as the primary indicator of joint attention. In contrast, SLP3 viewed engagement as the core marker, paying particular attention to whether the child demonstrated emotional or attentional alignment with the adult, even in the absence of explicit cues like gaze or pointing.
All three experts agreed that joint attention can manifest differently across developmental stages, influenced by a child’s evolving language, motor, and attention capacities. For younger children–particularly those under age two–they allowed for broader interpretations, including reliance on gestures, posture, or affective responses. For older children, they expected more deliberate and referential signals, such as consistent gaze shifts, pointing, or verbal referencing, reflecting increased communicative intention and control.
Building on these conceptual insights, we next asked each SLP to annotate segments from our curated video dataset. This annotation study allowed us to observe how their stated definitions translated into actual evaluations of strong and poor joint attention in real-world parent-child interactions.
3.6. Annotation Study Results: SLPs’ Observational Judgements of Joint Attention
During the annotation process, we observed that all three SLPs consistently relied on three core behavioural dimensions when evaluating joint attention. These dimensions–gaze, action, and vocalisation–served as the primary basis for assessing both the child’s engagement and the parent’s responsiveness. These cues were also frequently referenced in the SLPs’ verbal justifications, highlighting their prominence in expert reasoning. Below, we describe how each dimension was used in practice:
  • Gaze – Refers to where the child is looking, such as toward the parent’s face, hands, a shared object, or away from the interaction. Gaze cues are critical for interpreting visual attention and shared focus.
  • Action – Encompasses the child’s physical behaviours, including pointing, reaching, standing up, dragging objects, or attempting to disengage. These actions signal communicative intent and social participation.
  • Vocalisation – Includes all forms of vocal output, from babbling and laughter to verbal responses. Vocal cues provide insight into the child’s attempt to share attention or respond within the interaction.
This triadic lens emerged organically across SLPs and provided a shared framework for interpreting joint attention episodes in a structured yet flexible manner.
To facilitate structured analysis of expert annotations, we divided each video into uniform 5-second segments. While SLPs initially labelled joint attention quality by selecting start and end timestamps based on their judgement, we mapped these annotations onto the fixed segments, assigning each 5-second unit a corresponding label where overlap occurred. This mapping enabled us to translate variable-length judgements into a standardised format, making it easier to compare agreement across experts and analyse annotation patterns consistently across the dataset. Given that the activities–such as guided play or task-based exchanges–are generally short and focused, 5-second segments provided a practical and interpretable unit of analysis.
Table 3 summarises the distribution of joint attention labels assigned by each SLP, as well as the aggregated results based on majority agreement (i.e., at least two SLPs assigning the same label). Across all three experts, the majority of segments were labelled as Moderate, suggesting that many observed behaviours fell into an intermediate range that did not clearly indicate strong or poor joint attention. Notably, Poor labels were used relatively sparingly by SLP1 and only slightly more frequently by SLP2 and SLP3–reflecting a general reluctance to assign negative evaluations unless disengagement was highly evident.
The combined agreement row reveals that only 22.1% of segments were jointly considered Strong, and just 1.6% were jointly labelled as Poor, while over three-quarters (76.3%) were consistently judged as Moderate. This distribution highlights the interpretive nature of joint attention assessment and the need for careful alignment when designing systems to simulate expert reasoning.
Table 3. Distribution of Joint Attention labels across individual SLPs and the combined agreement set ($\geq$2 SLPs). Each SLP annotated all 638 five-second video segments. The combined agreement set excludes segments without at least two SLPs agreeing on the same label, resulting in 615 segments (23 segments removed due to lack of consensus).
ID Total Strong (n,%) Moderate (n,%) Poor (n,%)
SLP1 638 155 (24.3%) 472 (74.0%) 11 (1.7%)
SLP2 638 195 (30.6%) 392 (61.4%) 51 (8.0%)
SLP3 638 150 (23.5%) 437 (68.5%) 51 (8.0%)
Combined ($\geq$2 agree) 615 136 (22.1%) 469 (76.3%) 10 (1.6%)
3.6.1. Illustrative Examples of Consensus Among SLPs
While expert disagreement was common across many ambiguous segments, a smaller set of segments achieved full agreement among all three SLPs. These moments of consensus reveal the behavioural patterns that experts collectively interpret as clear evidence for strong, moderate, or poor joint attention. In general, jointly labelled Strong segments featured multiple forms of engagement–such as mutual gaze, verbal responsiveness, and shared task focus–while Poor segments were marked by disengagement, lack of social referencing, or solitary behaviour. Meanwhile, segments labelled Moderate often involved task participation without clear signs of shared social coordination. The following examples help clarify the implicit thresholds that delineate high-confidence assessments in expert judgement:
  • Video 19 Segment 001: In this segment, the parent placed coins on the table while explaining the setup (“So I’m going to make two rows of quarters”) and directed their gaze to the coins. The child, seated with hands clasped on the table, briefly vocalised (“oh”) and followed the parent’s actions with their gaze. The combination of attentiveness, vocal response, and shared focus led all three SLPs to rate the segment as Strong.
  • Video 2 Segment 004: In this segment, the parent sat on the floor, picked up Lego blocks, and labelled them by saying “Tower.” The child, lying on the floor, manipulated the blocks and verbally speculated about their play (“Oh, it could be a…”). Both parent and child kept their gaze on the Lego rather than each other, resulting in limited reciprocity. All three SLPs rated the segment as Moderate.
  • Video 9 Segment 001: In this segment, the parent sat on the floor facing the child and offered praise (“Ooh, I like the way you’re connecting those guys”) while looking toward the child and the toys. The child, kneeling on the floor and connecting toy pieces, kept their gaze down on the toys and did not respond verbally. The lack of reciprocity or shared focus led all three SLPs to rate the segment as Poor.
These consensus segments demonstrate how, despite individual stylistic differences, experts converge when multiple cues co-occur (for Strong), are present but non-social (for Moderate), or are entirely absent (for Poor). Such segments serve as useful anchors when training or evaluating AI models that aim to simulate SLP-like assessments.
3.6.2. Illustrative Examples of Diverging SLP Judgements
Although all three SLPs shared a general understanding of joint attention as a coordinated social process, their specific criteria and interpretive emphasis varied. These differences became especially evident when they encountered ambiguous or intermediate segments.
SLP1 adopted a balanced and multimodal approach. While gaze was still considered, it was treated as one of several contributing cues–alongside gestures, verbal referencing, emotional affect, and body orientation. SLP2 often acknowledged child-led play and nonverbal forms of engagement, showing flexibility in how joint attention could be expressed. Segments with inconsistent gaze but strong affective presence were still given high ratings when the child’s overall behaviour reflected social coordination.
SLP2, in contrast, placed strong emphasis on visual engagement, particularly gaze alternation between the adult and shared object. Eye contact and sustained visual referencing were seen as essential for identifying strong joint attention. In the absence of clear gaze cues, even segments involving verbal speech or physical interaction were often rated as moderate or poor. For example, SLP1 rated several segments as poor when the child appeared engaged but did not meet the baseline requirement of mutual gaze.
SLP3 took a functional and context-sensitive perspective, often evaluating segments based on the child’s communicative intent rather than adherence to typical behavioural markers. Rather than expecting conventional cues like pointing or verbalisation, SLP3 recognised alternative forms of participation–such as pushing away the parent’s hand or physically repositioning themselves–as valid signs of engagement. This expert was less concerned with surface-level eye contact and more attentive to whether the child demonstrated awareness and responsiveness in their own way.
To illustrate these differences, we highlight several segments where the three SLPs disagreed in their assessments. These examples underscore the unique interpretive lenses each expert applied:
  • Video 11 Segment 002: In this segment, the child sat on the floor playing with a toy airplane and said, “I want to go to the Bahamas. Fly there now.” While there was no gaze toward the parent, the child’s speech was contextually rich. SLP1, recognising the imaginative verbal output and engagement, rated it as Moderate. SLP2 labelled the segment as Poor due to the lack of visual coordination. SLP3 rated it Strong, interpreting the child’s verbal and motor actions as clear evidence of communicative intent.
  • Video 21 Segment 040: The child alternated gaze between a toy and the parent’s hands while reaching for the toy, but did not speak. SLP1 gave a Moderate rating, due to the absence of vocal engagement. SLP2 labelled this as Strong, citing the clear visual coordination. SLP3 judged it as Poor, arguing that the interaction lacked reciprocal cues or intentional signalling.
  • Video 21 Segment 017: In this segment, the child pointed to a toy lion without speaking. SLP1 rated it Moderate, perhaps noting the absence of emotional or verbal expression. SLP2 rated the interaction as Strong, identifying the combination of gaze and pointing as sufficient for joint attention. SLP3 gave it a Poor, interpreting the gesture as mechanical rather than socially directed.
Together, these examples demonstrate how gaze, vocalisation, and action were weighted differently by each expert, and how their background and theoretical orientation shaped their interpretation
4. Exploring Aligning MLLMs with SLPs: Behaviour Observations to Judgement
Building on our findings from the interview and annotation studies, we sought to model SLPs’ reasoning processes computationally. We designed a two-stage system that enables MLLMs to simulate how SLPs observe and judge joint attention. The first stage focuses on structured behavioural observation–using the three expert-derived dimensions of gaze, action, and vocalisation to prompt the MLLM to describe what is happening in each video segment. The second stage then assesses whether these structured descriptions improve the MLLM’s ability to judge the quality of joint attention. We evaluate this by comparing zero-shot and many-shot prompting strategies, and contrasting reasoning-based and non-reasoning model variants. Through this pipeline, we examine how well MLLMs can align with expert criteria in both perception and judgement.
4.1. Stage 1: Behaviour Description through Expert-Aligned Prompting
4.1.1. Method
With the rapid advancement of MLLMs, their application to human behaviour analysis presents a promising frontier. In this study, we adopt Gemini 2.5 Pro [^5] https://blog.google/products/gemini/gemini-2-5-pro-updates/ for its strong multimodal reasoning capabilities (yue2024mmmu; team2024gemini; fu2024video). However, prior research (shi2025towards) shows that directly prompting MLLMs to identify key interaction segments–especially for complex, socially embedded tasks like joint attention–often fails. Models continue to struggle with long video understanding, temporal grounding, and fine-grained cues such as gaze interpretation.
To address these limitations, we build on insights from our SLP interviews, which revealed two important features of joint attention evaluation: (1) key behavioural changes are rapid and often context-independent, and (2) expert assessments consistently rely on three observable cues–gaze, action, and vocalisation (see Table 2).
Guided by these principles, we divided each video into uniform 5-second segments and designed a structured prompting scheme to elicit behavioural summaries from the MLLM. For each segment, the model was asked to describe both the parent and the child’s behaviour along the three core dimensions. To encourage accurate, grounded behavioural descriptions, we designed a zero-shot prompt that guides the model to produce short, factual observations in natural language. By enforcing a consistent subject-verb-object structure, the prompt reflects the interpretive style of human annotators and reduces the likelihood of hallucinated or overly abstract outputs. The full prompt is shown below:
Structured Behaviour Description Prompt (Gaze-Action-Vocalisation):
You are watching a video of a parent interacting with a child.
For each participant (parent and child), describe their behaviour in three parts:
1. Gaze: Describe what or whom the person was looking at, using a natural language sentence with a subject-verb-object structure. Examples: - The child looked at the parent’s face. - The parent shifted gaze between the child and the toy. - The child stared at the blocks. - The parent looked away for a moment.
2. Action: Describe what the person physically did, using a short natural language sentence with a subject-verb-object structure. Examples: - The parent pointed at the red ball. - The child reached towards the puzzle pieces. - The parent lifted the toy truck. - The child clapped their hands.
3. Vocalisation: Transcribe or paraphrase what the person said, or describe any vocalisations using a short sentence with a subject-verb-object structure. Write ‘None’ if there were no vocalisations.
4.1.2. Result
We ran the full set of behaviour description prompts on an Ubuntu 22.04 LTS server via the official Gemini 2.5 Pro API, collecting MLLM-generated outputs for all 5-second video segments in our dataset.
To evaluate the MLLM’s behavioural description performance, we compared its outputs with reference annotations created by our research team. These reference annotations were constructed in two stages, following the same gaze-action-vocalisation structure used in the prompt. Importantly, our annotation process was informed by the interview findings with SLPs–particularly their attention to detail and tendency to distinguish subtle variations in gaze direction, gesture context, and vocalisation specificity.
For each segment, we began by manually reviewing the MLLM-generated outputs. In the first stage, we corrected instances where the generated description contradicted the observed behaviour–for example, changing “the child looked at the parent” to “the child looked down at the table” if no eye contact was actually made. In the second stage, we refined underspecified descriptions by aligning them with SLP-like granularity–for instance, revising “the child looked at the parent” to “the child looked at the parent’s hand” when the visual target was more specific. These two-step corrections ensured that reference labels reflected both factual accuracy and expert-informed attentional focus.
Based on this corrected reference set, we computed segment-level accuracy scores for each of the three behavioural fields. As shown in Table 4, the MLLM performed best in the action category, with a mean accuracy of 0.88, followed by vocalisation (0.87) and gaze (0.86). All three fields achieved perfect accuracy in at least one video, but gaze also had the lowest minimum, highlighting its difficulty for the model.
Table 4. Accuracy statistics by behavioural field
Field Mean Median Max Min
Action 0.8774 0.9464 1.0000 0.6250
Vocalisation 0.8708 0.9259 1.0000 0.5000
Gaze 0.8556 0.8750 1.0000 0.5000
To better understand model failure cases, we conducted a qualitative review of low-performing videos across the three behavioural fields. Several recurring issues emerged:
  • Speech role confusion was a major factor affecting vocalisation accuracy. For instance, in Video 24 and Video 25, the model consistently attributed child-like vocalisations to the parent-particularly when the parent mimicked the child’s babbling. Similar confusion was observed in Video 11, where rapid turn-taking and overlapping speech made speaker attribution unreliable.
  • Gaze misinterpretation often occurred when faces were partially occluded, or when the child looked at non-face targets such as hands or objects. In Video 11, the child’s gaze was repeatedly marked as “looking at the parent” despite clear visual evidence that the child was focused on the toy airplane.
  • Action detection errors were more frequent in unstructured scenes. For example, in Video 4, the child’s aggressive toy-hitting was not recognised, and in Video 15, a clear tray-passing motion was entirely missed. In Video 5, task misunderstanding led to mismatched labels (e.g., writing mistaken for eating).
These results demonstrate that our structured prompting approach substantially improved the reliability of MLLM-generated behavioural annotations, enabling the model to produce interpretable outputs that aligned with expert-labelled segment judgments in many cases. Compared to Shi et al. (shi2025towards), where average behavioural and gaze accuracy was around 62%, our structured prompting approach led to substantial improvements in both description accuracy and eye-contact interpretation.
4.2. Stage 2: Challenges in Simulating SLP Joint Attention Judgement with Prompting
In Stage 2, we investigated whether MLLMs could approximate SLPs’ judgements of joint attention when provided with the behavioural cues identified in Stage 1. This step shifts from describing observable behaviours to making evaluative, subjective judgements–a task that is inherently more challenging. We conducted an experiment with GPT-4.1, comparing zero-shot and many-shot prompting strategies, and evaluated model outputs against SLP annotations using precision, recall, and F1.
4.2.1. Dataset
We assembled a text-based agreement dataset from 25 parent–child interaction (PCI) videos analysed in Stage 1. In this stage, each segment was coded along three behavioural cue dimensions–gaze, action, and vocalisation–for both the parent and the child using AI. We reviewed the Stage 1 AI annotations of behaviour, corrected erroneous information, and consolidated them into $(\textit{parent, child}),\textit{label}$ pairs. Only segments with at least two SLPs in agreement were retained, and all cases of complete disagreement were removed. This resulted in a total of 615 agreed pairs.
4.2.2. Experiments
We employed GPT-4.1 (gpt-4.1-2025-04-14) [^6] https://platform.openai.com/docs/models/gpt-4.1–at present OpenAI’s most capable non-reasoning model–as the backbone model for text-based judgements. Evaluation was conducted on a video-by-video basis, where all segments from the same video were included together in the prompt to preserve contextual continuity. Each segment was represented as a structured observation in the form of paired $(\textit{parent},\textit{child})$ descriptions across gaze, action, and vocalisation, and the model was required to assign one overall label: Strong, Moderate, or Poor.
To assess how SLPs’ professional knowledge supports the alignment of MLLMs with expert judgement, we designed two prompting conditions: a zero-shot condition without access to SLP examples, and a many-shot condition that incorporated many SLP-annotated examples. The model was then asked to judge all segments of the held-out target video (grouped together to preserve contextual continuity), returning one label per segment in the prescribed format.
Drawing from our earlier interviews, we observed that SLPs typically arrive at a judgement through three steps: observing behavioural cues, reasoning about social coordination, and mapping these to categorical Judgements (e.g., strong, moderate, poor). We therefore designed the prompt structure to mirror this expert reasoning pipeline.
In the zero-shot condition, the prompt contained only task instructions, and the model produced judgements directly from the observational descriptions.
Zero-shot prompt:
You are a speech-language pathologist. Joint attention in a child refers to the ability to share attention with another person--typically the parent--by coordinating behaviours such as actions, vocalisations (e.g., speaking or making sounds), or gaze (looking at shared objects or people).
Please evaluate the quality of the child’s joint attention in each segment below based on their behaviours. Respond using the following format:
Segment 1: [Strong/Moderate/Poor]
Segment 2: ...
In the many-shot condition, we adopted a leave-one-video-out strategy (zhao2024learning): the held-out video served as the evaluation set, while annotated examples from all remaining videos-each consisting of $(\textit{parent},\textit{child})$ observations paired with a gold label-were embedded in the prompt as demonstrations. This design ensured that the model had access to the full range of SLP-provided knowledge from other videos when making judgements on the target segments. The procedure was repeated across all videos, so that every video was evaluated once as the hold-out, and the union of predictions from all folds yielded the complete set of model-generated judgement labels.
Below is a simplified illustration of the many-shot prompting structure used in our experiments.
Many-shot examples:
Below are labelled examples from other videos by real speech-language therapists. Each example shows a structured observational description and its judgement label.
Example 124
Parent:
- Action: The parent sat at the table with their hands clasped.
- Vocalisation: The parent said ‘‘Maybe, possibly’’.
- Gaze: The parent looked at the child and the game board.
Child:
- Action: The child placed a white game piece onto the game board.
- Vocalisation: None
- Gaze: The child looked at the game board and the pieces.
Judgement: Moderate
Example 178 Parent:
- Action: The parent sat at the table with her hands resting on it.
- Vocalisation: The parent said ‘‘Tricky. Ooh, you switched it. I like this.’’
- Gaze: The parent looked at the child and the game board.
Child:
- Action: The child moved a white game piece across the board.
- Vocalisation: None
- Gaze: The child looked down at the game board.
Judgement: Strong
Example 273 Parent:
- Action: The parent sat at the table with hands clasped.
- Vocalisation: The parent said ‘‘Are you getting the trains ready?’’
- Gaze: The parent looked at the child.
Child: - Action: The child pushed wooden train track pieces together on the table.
- Vocalisation: The child said ‘‘Train’’.
- Gaze: The child looked down at the train track pieces.
Judgement: Poor
4.2.3. Evaluation metrics
To assess overall alignment, we compared model-generated labels with the agreement label for each segment (i.e., the consensus label retained in the agreement dataset, requiring at least two SLPs to agree). We report accuracy; macro-precision, macro-recall, and macro-F1 (unweighted averages over Strong, Moderate, Poor to mitigate class imbalance); and Cohen’s $\kappa$ to quantify agreement beyond chance. All metrics are computed at the segment level across the full agreement dataset. In addition to macro-level scores, we also report per-class precision, recall, and F1 to characterise performance on each judgement category.
4.2.4. Result: Many-shot Prompting Strategy Yields Better Alignment
Overall performance
Many-shot prompting led to stronger alignment with expert consensus than zero-shot prompting (see Figure 5). Accuracy rose from 0.44 to 0.57, and Cohen’s $\kappa$ nearly doubled (0.10 → 0.18), indicating more consistent agreement beyond chance. Macro-precision also improved slightly (0.40 → 0.42), suggesting fewer false positives, while macro-recall decreased (0.53 → 0.46), implying a more conservative stance that missed some expert-labelled cases. Despite this trade-off, the overall macro-F1 increased from 0.34 to 0.40, reflecting a net gain in balanced performance.
Per-class performance: Strong and Moderate improve, Poor remains unreliable.
At the category level, many-shot prompting yields improvements for Strong and Moderate cases but not for Poor. For Strong, recall rises (0.48→0.60) and F1 improves (0.41→0.49), indicating more reliable detection of clear joint attention. For Moderate, which dominates the dataset, both precision (0.80→0.83) and recall (0.42→0.57) increase, producing the largest F1 gain (0.55→0.67). By contrast, Poor segments remain highly challenging: recall drops (0.70→0.20) and F1 declines (0.07→0.04). This reflects two issues. First, zero-shot prompting severely over-predicted poor cases, inflating recall but with almost no precision (0.04). many-shot prompting suppresses this over-prediction, yielding more conservative outputs but still very low precision (0.02). Second, the data distribution itself is highly imbalanced: across the dataset, Moderate accounts for 469 segments (73.5%), Strong for 136 (21.3%), while Poor is represented by only 10 examples (1.6%). Such scarcity makes it difficult to validate or learn reliable patterns for low-fidelity behaviours in a many-shot condition.
Figure 5. Radar plots comparing zero-shot (red) and many-shot (blue) prompting across overall and per-class evaluation metrics. many-shot consistently improves accuracy, macro-F1, and Cohen’s $\kappa$ in the overall condition, while also boosting performance on Strong and Moderate categories. Performance on the Poor category remains weak under both conditions, reflecting its limited representation in the dataset.:
5. Discussion
5.1. Greater Alignment at the Level of Behavioural Observations
SLPs showed strong alignment with one another with behavioural observation, consistently pointing to gaze, action, and vocalisation as the central dimensions for describing joint attention. Our structured prompting method leveraged this shared consensus and achieved similarly high alignment: MLLMs could reliably capture and reproduce these behavioural cues.
Because expert criteria were already aligned, observation proved to be a tractable entry point for human–AI alignment. By guiding MLLMs to “see” behaviours through these shared dimensions, we can achieve stable and reproducible outputs. This suggests that systems for analysing Parent-Child Interaction should prioritise clear, structured observation before attempting higher-level reasoning.
The relative ease of alignment in observation highlights a broader design lesson: when experts already share common descriptors, MLLMs can be scaffolded to replicate them reliably. Multimodal models are particularly well-suited to capture low-level, directly perceivable cues—for instance, gaze direction, vocal onset, or gesture trajectories—because such behaviours are already described consistently across experts. Achieving alignment in these cases mainly requires iterative confirmation with human experts, rather than inventing new criteria.
For example, in education (xing2024survey), teachers can consistently agree on whether a child raised a hand, spoke out loud, or looked at the teacher. In UX evaluation (vermeeren2010user), experts reliably agree on whether a user paused, clicked, or hovered over an interface element. In creative work critiques (alabood2023systematic), reviewers may all notice surface-level features such as colour saturation or the use of bold lines. These cues are easier for multimodal models to observe, making observation alignment relatively tractable.
5.2. Partial Alignment at the Level of Judgement
Overall alignment scores at the judgement stage were modest, reflecting the difficulty of capturing interpretive nuances. Nevertheless, we found that many-shot prompting consistently outperformed zero-shot, suggesting that incorporating examples grounded in SLP expertise helps models approximate expert reasoning more closely. This indicates that expert knowledge does play a role in supporting judgement-level alignment, even if the gains are incremental.
The results highlight the need for more deliberate strategies to improve judgement-level alignment. Few-shot prompting offers one pathway, but more sustained approaches such as fine-tuning with expert-labelled data or retrieval-augmented generation (RAG) may be necessary to embed richer interpretive cues. At the same time, because judgement involves a degree of subjectivity, future systems may need to actively explore how different forms of expertise shape interpretation, and provide mechanisms to surface or adapt to these perspectives rather than converge on a single fixed standard.
This raises a broader question about how to handle inherently subjective layers of expertise in human-AI alignment. At the observational level, alignment is more attainable because experts tend to converge on concrete cues such as gaze, action, or vocalisation. At the judgement level, however, interpretations vary and are often context-dependent, making consensus elusive. This challenge is not unique to joint attention: expert decision-making in many fields is notoriously tacit and difficult to distil into explicit rules, drawing on years of situated practice and subtle interpretive skill. For instance, in medicine, radiologists may differ in judging whether a faint shadow on a scan indicates pathology or normal variation. In law, judges interpreting similar facts may reach different conclusions based on contextual reasoning. In Parent-Child Interaction analysis, where interpretive diversity is intrinsic, the goal may be less about maximising consensus and more about enabling systems to represent, contrast, and adapt to multiple plausible viewpoints. This perspective opens up space for HCI and AI research to treat subjectivity not as noise to be eliminated, but as a critical design consideration in the development of judgment-sensitive systems.
In education, a raised hand may be judged by one teacher as evidence of engagement, while another views it as disruptive if out of turn. In UX evaluation, a pause may be seen by one expert as a usability breakdown, but by another as part of a normal learning curve. In creative work critiques, the same bold use of colour may be praised by one reviewer as innovative, yet criticised by another as unbalanced. These examples highlight that while observation alignment is attainable, judgement alignment may be less realistic or even undesirable. For exploratory domains like Parent-Child Interaction analysis, acknowledging and surfacing disagreement may be more valuable than enforcing consensus, as it allows users to reflect on alternative interpretations rather than rely on a single definitive output.
5.3. Limitations
Sample size
Our study involved only three SLPs, which necessarily limits the robustness and generalisability of the findings. While their insights were rich and directly informed our two-stage pipeline design, the small $N=3$ means the patterns we report should be read as illustrative rather than definitive. Future work will need to engage a larger and more diverse pool of practitioners to examine whether the alignment and misalignment patterns we observed hold across settings and populations.
Dataset characteristics
The dataset of 25 YouTube videos was skewed toward short, likely neurotypical interactions, introducing bias and constraining generalisability. Our contribution is exploratory, and we used demonstration-based recordings as a pragmatic choice given the difficulty of collecting authentic, neurodiverse, and naturally occurring Parent-Child Interaction data. However, this bias led to very few segments labelled as Poor, meaning our analysis primarily reflects alignment in the Strong category and provides only partial evidence of model–expert agreement overall. Notably, one SLP emphasised that detecting Poor joint attention is often clinically more critical, and suggested that recordings featuring neurodiverse children (e.g., autism-focused datasets) may not only be more representative of practice but also easier to annotate consistently. Expanding future datasets to include longer interactions, neurodiverse populations (benton2014diversity), and recordings from real-world home or school settings will be essential for improving ecological validity.
Positioning and Scope
It is important to clarify the scope of this work. We do not propose or validate any clinical framework for diagnosis or intervention, nor do we claim clinical applicability. Rather, we present an exploratory, case-based attempt to align model outputs with expert perspectives in parent–child interaction analysis. Our aim is to examine where alignment appears feasible at the level of observable behaviour, and where it becomes more challenging at the level of interpretive judgement, thereby surfacing design questions for future HCI systems. In this sense, the system is a probe rather than a contribution in itself: it is used to study alignment patterns, not to introduce a novel prompting technique or clinical tool. Any move toward practical use would require larger and more varied datasets, broader expert participation, and rigorous validation, which are beyond the scope of this study.
6. Conclusion
This paper presented an exploratory study of alignment in the context of parent–child interaction analysis. By comparing three SLP’ perspectives on joint attention with MLLM outputs, we identified a clear contrast: observation-level alignment was relatively robust, while judgement-level alignment remained elusive due to differences among experts themselves. These findings suggest that MLLMs can serve as reliable observers when scaffolded by shared behavioural cues, but that interpretive alignment requires more nuanced approaches and may benefit from exposing diverse perspectives rather than enforcing consensus.
Our contribution should be understood as a case-based investigation of alignment rather than a validated framework. The study highlights design opportunities for HCI and AI alignment research more broadly, where systems may act as collaborators that surface observations and alternative interpretations.
Appendix A Video Dataset Appendix
Youtube Videos (with link)
Behavioral Guidance and Skill Modeling
Video 2: Parent Models All PRIDE Skills & Active Ignoring- Parent-Child Interaction Therapy (PCIT)
Video 3: Therapist Coaches Big Ignore- Parent-Child Interaction Therapy (PCIT) for Older Children
Video 5: Parent Uses -Big Ignore- Technique- Parent-Child Interaction Therapy (PCIT) for Older Children
Video 6: Parent Models Skill to AVOID, Commands– Parent-Child Interaction Therapy (PCIT)
Video 8: Parent Models Skill to AVOID, Negative Talk– Parent-Child Interaction Therapy (PCIT)
Video 9: Parent Models PRIDE Skill -P for Praise– Parent-Child Interaction Therapy (PCIT)
Video 11: Parent Models PRIDE Skill -R for Reflection– Parent-Child Interaction Therapy (PCIT)
Video 12: Parent Models PRIDE Skill -I for Imitation– Parent-Child Interaction Therapy (PCIT)
Video 13: Parent Models PRIDE Skill -D for Description– Parent-Child Interaction Therapy (PCIT)
Video 14: Parent Models PRIDE Skill -E for Enthusiasm-Enjoyment– Parent-Child Interaction Therapy (PCIT)
Video 16: Parent Models Differential Attention-Active Ignoring- Parent-Child Interaction Therapy (PCIT)
Language and Cognitive Development
Video 17: A Typical 10-month-old on Piaget’s A-not-B task
Video 18: A Typical 3-year-old Sorting cards
Video 19: A typical child on Piaget’s conservation tasks
Video 20: ABA Sample Session (cards and chase)
Video 21: ABA Therapy - Learning about Animals
Video 22: ABA Therapy- Daniel - Communication
Video 23: Encouraging Language Development- A Positive Parent-Child Interaction
Video 24: Piaget - Object permanence failure (Sensorimotor Stage)
Video 25: Piaget - The A Not B Error (Sensorimotor Stage)
Daily Life Skills and Interaction
Video 1: Parent Models Skill to AVOID, Questions– Parent-Child Interaction Therapy (PCIT)
Video 4: Parent Models How to End Special Playtime- Parent-Child Interaction Therapy (PCIT)
Video 7: Parent & Child Check Sticker Chart- Parent-Child Interaction Therapy (PCIT) for Older Children
Video 10: Parent Meets CDI Goal Criteria- Parent-Child Interaction Therapy (PCIT) for Older Children
Video 15: Child Complies with Command in PDI- Parent-Child Interaction Therapy (PCIT) for Older Children

### v2

@@
 Figure 1.
-Conceptual overview: We investigate how MLLMs can be guided to align with SLPs in parent–child interaction through two key stages—Observation, using structured prompts based on behavioural cues such as gaze, action, and vocalisation; and Judgement, informed by few-shot learning with expert-derived practices.:
+Conceptual overview: We investigate how MLLMs can be guided to align with SLPs in parent–child interaction through two key stages—Observation, using structured prompts based on behavioural cues such as gaze, action, and vocalisation; and Judgement, informed by few-shot learning with expert-derived practices.
+This diagram illustrates how AI understands parent-child interactions by aligning with Speech-Language Pathologists (SLPs) and using cues like gaze, action, and vocalization.
 1. Introduction
@@
 Yet these behaviours are often subtle and difficult for caregivers to identify without professional support.
-This challenge is consequential: approximately one in six children in the United States experiences at least one developmental delay (robinson2017cdc), with language delays among the most prevalent (sunderajan_speech_2019).
-At the same time, access to speech-language pathologists (SLPs)–professionals who assess children’s communicative development by attending to multimodal cues such as gaze, gesture, vocalisation, and context–remains limited, particularly in rural or underserved communities (o2005barriers).
+This challenge is consequential: approximately one in six children in the United States experiences at least one developmental delay (Robinson, 2017), with language delays among the most prevalent (Sunderajan and Kanhere, 2019a).
+At the same time, access to speech-language pathologists (SLPs)–professionals who assess children’s communicative development by attending to multimodal cues such as gaze, gesture, vocalisation, and context–remains limited, particularly in rural or underserved communities (O’Callaghan et al., 2005).
 As a result, expert developmental knowledge and care often remain out of reach in everyday caregiving contexts.
-Despite recent advances in artificial intelligence (AI) for family and education (yuan2024designing; sun2024exploring; fiani2024exploring; nikkhah2024family; currin2024opportunities; su2024hidden), existing systems primarily assist with creating therapy materials or scaffolding caregiver engagement, rather than emulating the observing and judging practices of SLPs themselves.
+Despite recent advances in artificial intelligence (AI) for family and education (Yuan et al., 2024; Sun et al., 2024; Fiani et al., 2024; Nikkhah et al., 2024; Currin et al., 2024; Su et al., 2024), existing systems primarily assist with creating therapy materials or scaffolding caregiver engagement, rather than emulating the observing and judging practices of SLPs themselves.
 For example, Lewis et al.
-(lewis2025exploring) investigated how SLPs interact with AI-generated materials for children from culturally and linguistically diverse backgrounds, and Dangol et al.(dangol2025want) investigated how AI might support parents in delivering therapy at home, highlighting emotional and logistical challenges.
+(Lewis et al., 2025) investigated how SLPs interact with AI-generated materials for children from culturally and linguistically diverse backgrounds, and Dangol et al.(Dangol et al., 2025) investigated how AI might support parents in delivering therapy at home, highlighting emotional and logistical challenges.
 Yet in both cases, the evaluative expertise of SLPs remained external to the AI systems.
@@
 We focus on joint attention as a representative case: a core developmental milestone that SLPs typically evaluate through holistic, experience-based interpretations of children’s gaze, actions, and vocalisations.
-While MLLMs excel at structured tasks such as video captioning or temporal grounding, they struggle with socially nuanced behaviours that lack clear ground truth.
+While MLLMs excel at structured tasks such as video captioning or temporal grounding (Lu et al., 2024), they struggle with socially nuanced behaviours that lack clear ground truth.
+At the same time, recent advances suggest that MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (Gandhi et al., 2023; Jain et al., 2024), and continues to be an active area of research.
+Their ability to adapt to task-specific criteria through in-context learning makes them a promising candidate for bridging raw behavioural signals with the structured interpretive practices used by SLPs.
 We therefore investigate how expert criteria can be translated into structured representations that allow MLLMs to more closely align with expert judgment.
@@
 Drawing on interviews and annotation analyses with SLPs, we show that their evaluations hinge on three interrelated cues: gaze, action, and vocalisation.
-  • We design and evaluate an exploratory MLLM system that aligns with speech-language pathologists’ approaches to joint attention assessment in two stages: (1) observing fine-grained behavioural cues from parent-child interaction videos using expert-informed prompting, achieving up to 85% accuracy across dimensions; and (2) evaluating interaction quality using only structured behavioural descriptions, reaching over 57% average accuracy compared to expert labels.
+  • We design and evaluate an exploratory MLLM system that aligns with speech-language pathologists’ approaches to joint attention assessment in two stages: (1) observing fine-grained behavioural cues from parent-child interaction videos using expert-informed prompting, achieving up to 85% accuracy across dimensions; and (2) evaluating interaction quality using only structured behavioural descriptions, reaching over 64% average accuracy compared to expert labels.
   • We curate a segment-level dataset with expert-labelled joint attention behaviours and derive a set of practical design guidelines for building MLLM-based systems that align SLP’s observing and judging process.
@@
 2.1. Language Delay and the Limits of Current Parent Support
-Language delay is one of the most common developmental concerns in early childhood, with prevalence estimates ranging from 2.3% to 19% among children aged 2 to 7 years (mclaughlin2011speech).
-Without timely intervention, children with language delay are at increased risk for later difficulties in reading, attention regulation, and social interaction (sunderajan2019speech).
+Language delay is one of the most common developmental concerns in early childhood, with prevalence estimates ranging from 2.3% to 19% among children aged 2 to 7 years (McLaughlin, 2011).
+Without timely intervention, children with language delay are at increased risk for later difficulties in reading, attention regulation, and social interaction (Sunderajan and Kanhere, 2019b).
 SLPs play a key role in identifying and supporting children with language delays through structured interventions.
-However, limited access to professional services–whether due to long wait times, high cost, or service shortages–has led to growing interest in empowering parents to support their children at home (lieneman2017parent; straitstimes2024earlyintervention).
-Parent training programmes such as It Takes Two to Talk, More Than Words, and DIR Floortime have shown success in encouraging behaviours that promote early communication and social development (pepper2004talk; sussman1999words; brock2014statewide; dir_floortime).
+However, limited access to professional services–whether due to long wait times, high cost, or service shortages–has led to growing interest in empowering parents to support their children at home (Lieneman et al., 2017; The Straits Times, 2024).
+Parent training programmes such as It Takes Two to Talk, More Than Words, and DIR Floortime have shown success in encouraging behaviours that promote early communication and social development (Pepper and Weitzman, 2004; Sussman, 1999; Brock et al., 2014; The Interdisciplinary Council on Development and Learning, 2025).
 These interventions typically guide caregivers to be more responsive and attuned during everyday interactions, and have demonstrated benefits in improving the quantity and quality of parent-child communication.
-Yet while such programmes can teach parents strategies to encourage communication, they rarely equip caregivers to evaluate the developmental effectiveness of those interactions (finke2009all; gadberry2011survey; ganz2013impacts).
+Yet while such programmes can teach parents strategies to encourage communication, they rarely equip caregivers to evaluate the developmental effectiveness of those interactions (Finke et al., 2009; Gadberry, 2011; Ganz et al., 2013).
 For many parents, identifying nuanced developmental milestones can be difficult without professional guidance.
-This is especially true for joint attention (joint attention), one of the earliest and most important indicators of social-communicative development (mundy2007individual; tomasello1986joint).
-Determining whether joint attention is strong, weak, or absent often requires the kind of interpretive expertise that SLPs develop over years of practice and is challenging to translate into formal rules (o2005barriers).
+This is especially true for joint attention (joint attention), one of the earliest and most important indicators of social-communicative development (Mundy et al., 2007; Tomasello and Farrar, 1986).
+Determining whether joint attention is strong, weak, or absent often requires the kind of interpretive expertise that SLPs develop over years of practice and is challenging to translate into formal rules (O’Callaghan et al., 2005).
 Families outside of clinical environments rarely benefit from such expert judgment.
@@
 For instance, Chan et al.
-(chan2017wakey) developed WAKEY, which aimed to smooth morning routines by prompting parents to record schedules and track phrase usage.
+(Chan et al., 2017) developed WAKEY, which aimed to smooth morning routines by prompting parents to record schedules and track phrase usage.
 While effective in structuring communication, WAKEY relied entirely on manual input, without the ability to capture richer contextual signals.
 Subsequent work shifted towards real-time speech analysis to provide parents with more immediate feedback.
-TalkBetter (hwang2014talkbetter) analysed turn-taking and generated tailored recommendations to scaffold language development, while TalkLIME (song2016talklime) extended this approach by offering live metrics such as utterance count, initiation ratio, and turn-taking frequency.
+TalkBetter (Hwang et al., 2014) analysed turn-taking and generated tailored recommendations to scaffold language development, while TalkLIME (Song et al., 2016) extended this approach by offering live metrics such as utterance count, initiation ratio, and turn-taking frequency.
 More recent systems incorporated multimodal sensing to deepen their understanding of interaction dynamics.
 Captivate!
-(kwon2022captivate) leveraged gaze estimation and speech recognition to detect joint attention and suggest parent prompts during play, and AACessTalk (choi2025aacess) supported minimally verbal autistic children by combining vocabulary recommendations with contextual guidance, ultimately improving turn-taking and parental reflection over a two-week deployment.
+(Kwon et al., 2022) leveraged gaze estimation and speech recognition to detect joint attention and suggest parent prompts during play, and AACessTalk (Choi et al., 2025) supported minimally verbal autistic children by combining vocabulary recommendations with contextual guidance, ultimately improving turn-taking and parental reflection over a two-week deployment.
 Parallel to these system developments, researchers have also explored how parents can be supported in adopting therapy strategies more effectively.
 Dangol et al.
-(dangol2025want) investigated how AI might address emotional and logistical barriers to at-home therapy, while Li et al.
-(li2025asd) introduced ASD-HI, a multimodal dataset and baseline model designed to detect strategies and assess fidelity in parent-led interventions.
+(Dangol et al., 2025) investigated how AI might address emotional and logistical barriers to at-home therapy, while Li et al.
+(Li et al., 2025) introduced ASD-HI, a multimodal dataset and baseline model designed to detect strategies and assess fidelity in parent-led interventions.
 Together, this trajectory reflects a shift from manual, parent-driven tools to increasingly automated, multimodal, and expert-aligned systems, yet few have investigated whether AI can evaluate interaction quality against professional standards–a gap our work explores by aligning MLLM-based analysis with SLP judgement.
@@
 For example, Lewis et al.
-(lewis2025exploring) examined how SLPs interact with AI-generated visual content when working with children from culturally and linguistically diverse backgrounds.
+(Lewis et al., 2025) examined how SLPs interact with AI-generated visual content when working with children from culturally and linguistically diverse backgrounds.
 While their study highlighted important concerns regarding representation, bias, and contextual mismatch, it focused primarily on content usability.
 Their work does not address how AI might support or simulate the evaluative reasoning processes that SLPs use when observing and interpreting children’s behaviours in real time.
-With the rapid progress of MLLMs across tasks such as video captioning and temporal grounding (lu2024gpt), researchers have begun to explore their potential for analysing human behaviour–particularly in socially grounded contexts such as parent-child interaction.
-MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (gandhi2023multimodal; jain2024vcoder).
-Thanks to their in-context learning capabilities, MLLMs can flexibly adapt to new interaction settings with minimal data (dos2023composite; whitehead2024generative), making them promising for contexts that require domain-specific reasoning.
+With the rapid progress of MLLMs across tasks such as video captioning and temporal grounding (Lu et al., 2024), researchers have begun to explore their potential for analysing human behaviour–particularly in socially grounded contexts such as parent-child interaction.
+MLLMs can interpret complex verbal and non-verbal behaviour across video, audio, and text, and generate descriptive summaries of human activity (Gandhi et al., 2023; Jain et al., 2024).
+Thanks to their in-context learning capabilities, MLLMs can flexibly adapt to new interaction settings with minimal data (Dos Santos Melicio et al., 2023; Whitehead et al., 2024), making them promising for contexts that require domain-specific reasoning.
 Crucially, these models shift the output from fixed metrics or visualisation into natural language descriptions–potentially bridging the gap between raw behavioural signals and expert interpretation.
-Zheng et al.(zheng2024soap) introduced SOAP.AI, a system that enables experts to interact with MLLMs through in-context prompting and collaborative task design, and supports the automatic generation of initial behavioural segments–such as those relevant to joint attention–in domains like healthcare and education.
+Zheng et al.(Zheng et al., 2024) introduced SOAP.AI, a system that enables experts to interact with MLLMs through in-context prompting and collaborative task design, and supports the automatic generation of initial behavioural segments–such as those relevant to joint attention–in domains like healthcare and education.
 However, SOAP.AI was designed primarily as a tool for expert users, without evaluating the reliability of its outputs or aligning them with SLP judgement.
-Extending this line of inquiry, Zheng et al.(zheng2025ai) subsequently explored how AI-generated documentation could more directly support SLPs, identifying four opportunities and proposing three fluidity-focused design goals for future systems.
+Extending this line of inquiry, Zheng et al.(Zheng et al., 2025) subsequently explored how AI-generated documentation could more directly support SLPs, identifying four opportunities and proposing three fluidity-focused design goals for future systems.
 However, while SOAP.AI exemplifies how AI can augment expert workflows, it is designed primarily for expert users and does not evaluate the reliability of its outputs or align them with SLP judgement, leaving open the question of whether MLLMs can simulate expert evaluative capabilities.
-While these approaches centre on supporting SLPs in their clinical workflows, they do not address a critical question: can current technologies simulate the observational and interpretive reasoning that SLPs apply when evaluating parent-child interactions?
-Bridging this gap requires rethinking how AI systems can align with expert judgement–not just to assist SLPs, but to empower parents with meaningful, expert-level insights into their everyday interactions.
-Our work explores the foundational alignment of MLLMs with SLP reasoning, laying the basis for more reliable support to both experts and caregivers.
-2.4. Towards SLP-Aligned Observation and Judgement in Parent-Child Interaction
-The first attempt to align MLLMs with the evaluative judgement of SLPs was introduced by Shi et al.
-(shi2025towards), who prompted MLLMs with definitions of joint attention to directly identify strong or poor joint attention segments through a temporal grounding task.
-Their results showed poor performance, primarily due to the models’ limited sensitivity to fine-grained gaze cues–particularly eye contact, which is central to joint attention dynamics.
-These findings highlight the need for closer alignment between MLLM interpretation and the evaluative reasoning used by SLPs, rather than relying solely on definition-based prompting.
+While these approaches centre on supporting SLPs in their clinical workflows, they do not address a critical question: can current technologies simulate the observational and interpretive judging that SLPs apply when evaluating parent-child interactions?
+While general-purpose systems such as Personalized Judge (Dong et al., 2024) show that LLM-as-Judge can be effective in broad settings, their limitations become clear in expertise-heavy domains.
+Szymanski et al.
+(Szymanski et al., 2025; Chen et al., 2024) show that although LLMs align moderately with subject-matter experts in broad fields such as dietetics and mental health (64–68% agreement), their consistency drops on domain-specific questions that rely on professional judgement.
+Parent–child interaction illustrates this gap even more sharply: behaviours like joint attention require interpreting subtle, context-dependent cues rather than following fixed criteria.
+Prior attempts to prompt MLLMs for joint-attention grounding (Shi et al., 2025) yielded near-chance performance, largely due to the models’ insensitivity to fine-grained gaze signals.
+These patterns suggest that, despite general promise, LLM-as-Judge remains insufficient for specialised behavioural assessment without deeper alignment to expert reasoning.
 Our work continues this exploration by examining how MLLMs can be guided toward more SLP-aligned interpretations of parent-child interaction.
@@
 Accordingly, in the subsequent stage, we recruited SLPs as our human experts.
-Figure 2. Flow chart of the video screening process :
+Figure 2. Flow chart of the video screening process
+This diagram illustrates our video screening process, which began with a keyword search that yielded 198 videos.
+We then excluded videos based on our inclusion and exclusion criteria, resulting in a final set of 25 videos for the study.
 We curated a final dataset of 25 parent-child interaction videos (Video 1-25) spanning developmental stages from infancy to early school age (0-8 years), with most focusing on preschool years.
@@
 These videos were openly accessible and chosen with the intention of enhancing accessibility at this early stage.
-The videos in the dataset covers three categories: (1) behavioural guidance and skill modelling [^1] https://www.youtube.com/watch?v=YUkujhg6j6w (n=10), where parents demonstrate techniques such as PRIDE skills (masse2018taking) or “Big Ignore” (woodfield2021time); (2) language and cognitive development [^2] https://www.youtube.com/watch?v=rVqJacvywAQ (n=9), showing tasks like topic discussions or Piaget’s experiments; and (3) daily life skills and interaction [^3] https://www.youtube.com/watch?v=N3wAPLXd7I0 (n=6), including reward chart reviews or shared playtime.
+The videos in the dataset covers three categories: (1) behavioural guidance and skill modelling [^1] https://www.youtube.com/watch?v=YUkujhg6j6w (n=10), where parents demonstrate techniques such as PRIDE skills (Masse et al., 2018) or “Big Ignore” (Woodfield et al., 2021); (2) language and cognitive development [^2] https://www.youtube.com/watch?v=rVqJacvywAQ (n=9), showing tasks like topic discussions or Piaget’s experiments; and (3) daily life skills and interaction [^3] https://www.youtube.com/watch?v=N3wAPLXd7I0 (n=6), including reward chart reviews or shared playtime.
 Examples from each category are illustrated in Figure 3.
-(a) Behavioural Guidance and Skill Modelling: Parent uses the ”Big Ignore” (woodfield2021time) technique while the child continues painting (Video 3 (lieneman2023)).
-(b) Language and Cognitive Development: Parent and child play a toy-hiding game to foster engagement. (Video1 10 (Lieneman2024))
-(c) Daily Life Skills and Interaction: Parent and child build a marble run together, encouraging planning (Video2 (adam2013)).
-Figure 3. Examples from three categories in our dataset: behavioural guidance, language development, and daily life interaction.:
+(a) Behavioural Guidance and Skill Modelling: Parent uses the ”Big Ignore” (Woodfield et al., 2021) technique while the child continues painting (Video 3 (Lieneman, 2023a)).
+(b) Language and Cognitive Development: Parent and child play a toy-hiding game to foster engagement. (Video1 10 (Lieneman, 2023b))
+(c) Daily Life Skills and Interaction: Parent and child build a marble run together, encouraging planning (Video2 (Adam, 2013)).
+Figure 3. Examples from three categories in our dataset: behavioural guidance, language development, and daily life interaction.
+This diagram shows example videos from our dataset, categorized into three types: Behavioural Guidance and Skill Modelling (left), where a parent uses the ’Big Ignore’ technique while a child paints; Language and Cognitive Development (middle), where a parent and child play a toy-hiding game; and Daily Life Skills and Interaction (right), where they build a marble run together to encourage plannings
 3.2. Participants
 Our three expert SLPs were all female, aged between 30 and 59 years (see Table 1).
-All have extensive clinical experience working directly with children, particularly those on the autism spectrum, and have received training in a wide range of evidence-based intervention programmes, including Hanen’s It Takes Two to Talk (pepper2004talk), More Than Words and DIR Floortime (dir_floortime).
+All have extensive clinical experience working directly with children, particularly those on the autism spectrum, and have received training in a wide range of evidence-based intervention programmes, including Hanen’s It Takes Two to Talk (Pepper and Weitzman, 2004), More Than Words and DIR Floortime (The Interdisciplinary Council on Development and Learning, 2025).
 Their clinical practice spans public and private sectors and includes diverse settings such as early intervention centres, preschools, home-based care, and multidisciplinary teams.
@@
 Given their specialised training and long-standing experience, particularly with children on the autism spectrum, who often experience joint attention difficulties–the SLPs were well-qualified to identify and evaluate strong and poor instances of joint attention.
-Table 1. Profiles of the three expert SLPs who participated in this study, including their age group, academic qualifications, specialist certifications, and experience years.
 ID Age Group Qualifications Experience and Training
@@
 SLP3 55-59 BSc in Speech Pathology; Msc in Communication Disorders; Certified in Hanen (It Takes Two to Talk, More Than Words); DIR Floortime 23 years of paediatric therapy experience, primarily in school settings, with extensive international practice experience
+Table 1. Profiles of the three expert SLPs who participated in this study, including their age group, academic qualifications, specialist certifications, and experience years.
 3.3. Study Protocol
@@
 Each annotation included their short explanation justifying their decision and their suggestions for how Parent-Child Interaction could be supported.
-We used our custom-built video annotation tool to facilitate the SLPs’ judgement process (see Figure 4, shown here with a screenshot from Video 24444https://www.youtube.com/watch?v=rVqJacvywAQ (adam2013)).
+We used our custom-built video annotation tool to facilitate the SLPs’ judgement process (see Figure 4, shown here with a screenshot from Video 24444https://www.youtube.com/watch?v=rVqJacvywAQ (Adam, 2013)).
 Post-task interview (20 minutes)
@@
 Timeline Annotation Area for selecting and labelling segments as strong or poor joint attention; Note-Taking Area for recording justifications or observations; and Control Button Area for task submission and navigation.
-:
+This figure shows our video annotation tool, designed to support SLPs in their judgment process.
+It features four main components: a Video Playback Area (shown with a screenshot from Video 24) for watching interactions, a Timeline Annotation Area for labeling joint attention as ”strong” or ”poor,” a Note-Taking Area for recording observations, and a Control Button Area for task submission and navigation.
 3.4. Data Analysis
-The audio recordings were transcribed using WhisperX (bain2023whisperx) and manually verified.
+The audio recordings were transcribed using WhisperX (Bain et al., 2023) and manually verified.
 We analysed the qualitative data by combining the SLPs’ transcripts from the interviews with their annotations during the video tasks.
@@
 Combined ($\geq$2 agree) 615 136 (22.1%) 469 (76.3%) 10 (1.6%)
+To quantify the level of consistency among the three experts, we computed inter-annotator agreement using Krippendorff’s $\alpha$ with an ordinal scale, which reflects the ordered nature of the Strong, Moderate, and Poor labels.
+The overall $\alpha$ was 0.19, indicating low consistency among the three SLPs.
+This value indicates low consistency among the three SLPs.
+In combination with the heavily skewed label distribution, where each of the three experts marked more than 60% of segments as Moderate, the result suggests that disagreements often arose in borderline cases rather than in clearly distinct behaviours.
+Pairwise analyses further reveal asymmetries in expert alignment.
+The highest agreement occurred between SLP1 and SLP2 ($\alpha=0.31$), while substantially lower agreement was observed for SLP1–SLP3 ($\alpha=0.15$) and SLP2–SLP3 ($\alpha=0.11$).
+These differences suggest that the three SLPs do not form a single coherent “consensus group”; instead, their judgments cluster into at least two partially aligned perspectives.
+Taken together, the low overall $\alpha$ and uneven pairwise consistency indicate that expert interpretation of joint attention is inherently variable, especially in segments where behaviours fall between categories.
+This variability reinforces the importance of modelling expert diversity rather than assuming a single ground truth.
 3.6.1. Illustrative Examples of Consensus Among SLPs
@@
 While gaze was still considered, it was treated as one of several contributing cues–alongside gestures, verbal referencing, emotional affect, and body orientation.
-SLP2 often acknowledged child-led play and nonverbal forms of engagement, showing flexibility in how joint attention could be expressed.
+SLP1 often acknowledged child-led play and nonverbal forms of engagement, showing flexibility in how joint attention could be expressed.
 Segments with inconsistent gaze but strong affective presence were still given high ratings when the child’s overall behaviour reflected social coordination.
@@
 In the absence of clear gaze cues, even segments involving verbal speech or physical interaction were often rated as moderate or poor.
-For example, SLP1 rated several segments as poor when the child appeared engaged but did not meet the baseline requirement of mutual gaze.
+For example, SLP2 rated several segments as poor when the child appeared engaged but did not meet the baseline requirement of mutual gaze.
 SLP3 took a functional and context-sensitive perspective, often evaluating segments based on the child’s communicative intent rather than adherence to typical behavioural markers.
@@
 SLP3 gave it a Poor, interpreting the gesture as mechanical rather than socially directed.
-Together, these examples demonstrate how gaze, vocalisation, and action were weighted differently by each expert, and how their background and theoretical orientation shaped their interpretation
+Together, these examples demonstrate how gaze, vocalisation, and action were weighted differently by each expert, and how their background and theoretical orientation shaped their interpretation.
+This variability in interpretive emphasis directly aligns with the earlier statistical results, where overall inter-annotator agreement remained low ($\alpha=0.19$), reflecting the inherent subjectivity of these judgements.
 4. Exploring Aligning MLLMs with SLPs: Behaviour Observations to Judgement
@@
 With the rapid advancement of MLLMs, their application to human behaviour analysis presents a promising frontier.
-In this study, we adopt Gemini 2.5 Pro [^5] https://blog.google/products/gemini/gemini-2-5-pro-updates/ for its strong multimodal reasoning capabilities (yue2024mmmu; team2024gemini; fu2024video).
-However, prior research (shi2025towards) shows that directly prompting MLLMs to identify key interaction segments–especially for complex, socially embedded tasks like joint attention–often fails.
+In this study, we adopt Gemini 2.5 Pro [^5] https://blog.google/products/gemini/gemini-2-5-pro-updates/ for its strong multimodal reasoning capabilities (Yue et al., 2024; Team et al., 2024; Fu et al., 2025).
+A further practical consideration is that Gemini 2.5 Pro is, at present, the only widely accessible API model that accepts direct video inputs.
+Other foundation models, including the latest OpenAI offerings, currently support only image-based visual inputs, which makes them less suitable for segment-level behavioural captioning in continuous parent-child interaction videos.
+However, prior research (Shi et al., 2025) shows that directly prompting MLLMs to identify key interaction segments–especially for complex, socially embedded tasks like joint attention–often fails.
 Models continue to struggle with long video understanding, temporal grounding, and fine-grained cues such as gaze interpretation.
@@
 The full prompt is shown below:
-Structured Behaviour Description Prompt (Gaze-Action-Vocalisation):
+Structured Behaviour Description Prompt
+(Gaze-Action-Vocalisation)
 You are watching a video of a parent interacting with a child.
@@
 We ran the full set of behaviour description prompts on an Ubuntu 22.04 LTS server via the official Gemini 2.5 Pro API, collecting MLLM-generated outputs for all 5-second video segments in our dataset.
-To evaluate the MLLM’s behavioural description performance, we compared its outputs with reference annotations created by our research team.
-These reference annotations were constructed in two stages, following the same gaze-action-vocalisation structure used in the prompt.
-Importantly, our annotation process was informed by the interview findings with SLPs–particularly their attention to detail and tendency to distinguish subtle variations in gaze direction, gesture context, and vocalisation specificity.
-For each segment, we began by manually reviewing the MLLM-generated outputs.
-In the first stage, we corrected instances where the generated description contradicted the observed behaviour–for example, changing “the child looked at the parent” to “the child looked down at the table” if no eye contact was actually made.
-In the second stage, we refined underspecified descriptions by aligning them with SLP-like granularity–for instance, revising “the child looked at the parent” to “the child looked at the parent’s hand” when the visual target was more specific.
-These two-step corrections ensured that reference labels reflected both factual accuracy and expert-informed attentional focus.
-Based on this corrected reference set, we computed segment-level accuracy scores for each of the three behavioural fields.
-As shown in Table 4, the MLLM performed best in the action category, with a mean accuracy of 0.88, followed by vocalisation (0.87) and gaze (0.86).
-All three fields achieved perfect accuracy in at least one video, but gaze also had the lowest minimum, highlighting its difficulty for the model.
-Table 4. Accuracy statistics by behavioural field
+To evaluate the MLLM’s behavioural description performance, we first conducted a structured human review of all model-generated outputs for each segment.
+Two annotators from our research team independently examined every 5-second segment.
+For each segment, they reviewed all six model-generated entries (parent/child $\times$ gaze/action/vocalisation) and assigned a binary correctness label: 1 if the description accurately matched the observed behaviour in the video, and 0 otherwise.
+For each of the three behavioural fields, correctness was determined according to the subject–verb–object structure specified in the prompt.
+For gaze, an entry was marked as correct only if both the subject and the gaze target (object) were accurate, such as distinguishing between looking at the parent’s face or looking at the blocks held in the parent’s hand.
+For action, correctness required that the subject and the action described were consistent with the video, for example identifying whether the child was using a hand to pick up a block or performing another distinct motor behaviour.
+For vocalisation, an entry was considered correct only when the described verbal content matched what was actually spoken in the segment, including both who produced the speech and the approximate content of the utterance.
+These field-specific criteria ensured that the binary labels reflected the accuracy of the subject–verb–object components rather than surface-level wording.
+The two annotators achieved an inter-annotator agreement of Cohen’s $\kappa=0.92$, indicating high consistency.
+Any disagreements were resolved through discussion, and most arose from fine-grained distinctions in gaze descriptions.
+For example, the model sometimes output “child look at parent” when the behaviour should instead specify whether the child was looking at the parent’s face or hand.
+Based on these human-verified binary correctness labels, we computed accuracy for each behavioural field at the level of entire videos.
+For each video, we aggregated all segment-level binary labels for that field and calculated a per-video accuracy score.
+Table 4 reports the mean, median, maximum, and minimum of these per-video accuracies across the full dataset.
+As shown in the table, the MLLM performed best in the action category, with a mean per-video accuracy of 0.88, followed by vocalisation (0.87) and gaze (0.86).
+All three fields achieved perfect accuracy in at least one video, although gaze also had the lowest minimum accuracy, highlighting its difficulty for the model.
+Table 4. Per-video aggregated accuracy statistics. For each field, accuracies were computed per video, and the table summarises the mean, median, maximum, and minimum across all videos.
 Field Mean Median Max Min
@@
 These results demonstrate that our structured prompting approach substantially improved the reliability of MLLM-generated behavioural annotations, enabling the model to produce interpretable outputs that aligned with expert-labelled segment judgments in many cases.
-Compared to Shi et al.
-(shi2025towards), where average behavioural and gaze accuracy was around 62%, our structured prompting approach led to substantial improvements in both description accuracy and eye-contact interpretation.
+Shi et al.
+(Shi et al., 2025) first introduced this behavioural description task and dataset, reporting that the model achieved approximately 62% behavioural accuracy within the 23% of segments where it successfully detected gaze cues.
+Building on this prior setup, our structured prompting approach ensures that the model receives explicit action, vocalisation, and gaze information for every segment and thus produces complete descriptions for 100% of the dataset.
+Using a comparable combined accuracy metric that marks a segment as correct only when all three dimensions are correct, our method achieved 81% accuracy.
+This reflects a substantial improvement in both overall description quality and gaze-related interpretation.
 4.2. Stage 2: Challenges in Simulating SLP Joint Attention Judgement with Prompting
@@
 4.2.1. Dataset
-We assembled a text-based agreement dataset from 25 parent–child interaction (PCI) videos analysed in Stage 1.
-In this stage, each segment was coded along three behavioural cue dimensions–gaze, action, and vocalisation–for both the parent and the child using AI.
-We reviewed the Stage 1 AI annotations of behaviour, corrected erroneous information, and consolidated them into $(\textit{parent, child}),\textit{label}$ pairs.
-Only segments with at least two SLPs in agreement were retained, and all cases of complete disagreement were removed.
-This resulted in a total of 615 agreed pairs.
+The Stage 2 dataset was derived from the model-generated segment descriptions produced in Stage 1.
+For each 5-second segment across the 25 parent–child interaction videos, the model output six behavioural entries (parent/child $\times$
+gaze/action/vocalisation).
+Two annotators independently evaluated the correctness of each entry based on the video, and any disagreements were resolved through discussion until consensus was reached.
+This consensus served as the ground truth for determining which descriptions required correction.
+We then applied targeted revisions only to the entries marked as incorrect during the agreement process.
+Following a minimal-editing principle, we retained the original subject–verb–object structure whenever possible and corrected only the behavioural element that was factually inaccurate or mis-specified (e.g., replacing “the child looked at the parent” with “the child looked down at the table” if no eye contact occurred).
+When the model’s description was underspecified but directionally correct, we refined it to align with SLP-like granularity, such as specifying the actual attentional target (e.g., “the child looked at the parent’s hand”).
+The resulting set of minimally corrected and consensus-verified descriptions constituted the input to the Stage 2 judging process.
+[^6] We will release the anonymised annotated dataset and analysis materials upon acceptance to support transparency and reproducibility.
 4.2.2. Experiments
-We employed GPT-4.1 (gpt-4.1-2025-04-14) [^6] https://platform.openai.com/docs/models/gpt-4.1–at present OpenAI’s most capable non-reasoning model–as the backbone model for text-based judgements.
-Evaluation was conducted on a video-by-video basis, where all segments from the same video were included together in the prompt to preserve contextual continuity.
-Each segment was represented as a structured observation in the form of paired $(\textit{parent},\textit{child})$ descriptions across gaze, action, and vocalisation, and the model was required to assign one overall label: Strong, Moderate, or Poor.
+For the Stage 2 judging process, we used GPT-4.1 (gpt-4.1-2025-04-14) [^7] https://platform.openai.com/docs/models/gpt-4.1 as the backbone model for all text-based decisions.
+Stage 2 mirrors how SLPs conduct holistic assessment: instead of observing isolated 5-second segments, the model must read the entire sequence of segment-level descriptions from a full video and determine which portions correspond to Strong, Moderate, or Poor joint attention.
+This requires integrating behavioural evidence distributed across the interaction, similar to how an SLP reviews the full temporal flow of gaze, action, and vocalisation before forming a judgement.
+The task therefore primarily requires stable long-context understanding, enabling the model to integrate behavioural descriptions that span the full interaction rather than treating segments independently.
+Although no existing LLM-as-judge benchmark directly evaluates this form of segment-level behavioural scoring, the Stage 2 task imposes requirements that differ substantially from Stage 1.
+Stage 1 relies on Gemini 2.5 Pro, a strong video–reasoning model that excels at extracting behavioural cues from short, self-contained 5-second clips.
+In contrast, Stage 2 mirrors how SLPs make holistic judgements: the model must read the entire sequence of segment descriptions for a full video, integrate behavioural evidence distributed across many segments, and determine which portions correspond to Strong, Moderate, or Poor joint attention.
+In our dataset, the longest 12-minute video produced approximately 60,000 tokens of segment-level descriptions, meaning the Stage 2 model must process and integrate an input roughly of this size for a single judgement.
+As a result, long-context reliability, rather than step-by-step reasoning, becomes the primary requirement for Stage 2.
+At the time of our experiments, GPT-4.1 was the strongest publicly available model for long-context understanding.
+It supports a verified 1M-token context window and demonstrates superior performance compared with GPT-4o and o3 models on long-context retrieval and multi-hop long-context reasoning [^8] https://openai.com/index/gpt-4-1/.
+These benchmarks indicate that GPT-4.1 is particularly reliable at retrieving, tracking, and integrating information scattered across extended contexts.
+Such characteristics align directly with the needs of Stage 2, where the model must synthesise behavioural descriptions across an entire video without speculative inference.
+For these reasons, we selected GPT-4.1 as the backbone for Stage 2 rather than reusing the Stage 1 model or employing a specialised reasoning model.
+Evaluation in Stage 2 followed the same holistic procedure used by SLPs when annotating parent–child interactions.
+All segments from the same video were provided together in a single prompt, allowing the model to review the full sequence of behaviours before making any judgement.
+Each segment was represented as a structured pair of $(\textit{parent},\textit{child})$ observations across gaze, action, and vocalisation.
+Rather than assigning labels to segments independently, the model was required to integrate information across the entire video and identify which parts of the interaction corresponded to Strong, Moderate, or Poor joint attention, mirroring the way SLPs derive their final ratings from the full interaction.
 To assess how SLPs’ professional knowledge supports the alignment of MLLMs with expert judgement, we designed two prompting conditions: a zero-shot condition without access to SLP examples, and a many-shot condition that incorporated many SLP-annotated examples.
@@
 Segment 2: ...
-In the many-shot condition, we adopted a leave-one-video-out strategy (zhao2024learning): the held-out video served as the evaluation set, while annotated examples from all remaining videos-each consisting of $(\textit{parent},\textit{child})$ observations paired with a gold label-were embedded in the prompt as demonstrations.
+In the many-shot condition, we adopted a leave-one-video-out strategy (Zhao et al., 2024): the held-out video served as the evaluation set, while annotated examples from all remaining videos-each consisting of $(\textit{parent},\textit{child})$ observations paired with a gold label-were embedded in the prompt as demonstrations.
 This design ensured that the model had access to the full range of SLP-provided knowledge from other videos when making judgements on the target segments.
@@
 Judgement: Poor
+For both the zero-shot and many-shot settings, we evaluated the model using three decoding temperatures (0, 0.3, and 0.5).
+Because the judgment task requires stable and reliable decisions rather than highly diverse generations, we intentionally restricted the temperature range to relatively low values.
+To account for the stochasticity introduced at temperatures above zero, we conducted three independent runs for each non-deterministic setting (0.3 and 0.5).
+Temperature 0 is deterministic and therefore required only a single run.
 4.2.3. Evaluation metrics
@@
 4.2.4. Result: Many-shot Prompting Strategy Yields Better Alignment
+Temperature effects.
+Figure 5 shows clear temperature-dependent differences in model reliability across both prompting modes.
+Temperature 0.3 consistently produced the strongest and most stable performance: in the many-shot condition it achieved the highest accuracy ($0.64\pm 0.01$), macro-F1 ($0.44\pm 0.01$), and Cohen’s $\kappa$ ($0.26\pm 0.02$), while zero-shot prompting likewise reached its best scores at this setting (macro-F1 $0.351\pm 0.005$, $\kappa=0.103\pm 0.006$).
+In contrast, temperature 0 underperformed in both modes (e.g., many-shot macro-F1 =0.40=0.40, zero-shot macro-F1 =0.344=0.344).
+Temperature 0.5 yielded slightly lower means and substantially higher variance; for example, in many-shot prompting accuracy dropped to $0.61\pm 0.05$ and macro-F1 to $0.42\pm 0.02$, suggesting that excessive sampling randomness reduces stability.
+Overall, temperature 0.3 offers the best balance of accuracy and robustness, and we therefore adopt it as the default for subsequent analyses.
+Figure 5.
+Effect of decoding temperature on model reliability for zero-shot and many-shot prompting.
+Each subplot shows mean performance with standard-deviation error bars across temperatures 0, 0.3, and 0.5 for accuracy, macro-precision, macro-recall, macro-F1, and Cohen’s $\kappa$.
+Across all metrics, temperature 0.3 provides the most reliable overall performance, while temperature 0 underperforms and temperature 0.5 exhibits reduced stability due to increased sampling randomness.
+Many-shot prompting benefits more from moderate stochasticity than zero-shot prompting, which remains comparatively less sensitive to temperature changes.
+These results support the choice of temperature 0.3 for subsequent analyses.
+Five side-by-side line plots compare accuracy, macro-precision, macro-recall, macro-F1, and Cohen’s kappa across temperatures 0, 0.3, and 0.5 for zero-shot and many-shot prompting.
+Temperature 0.3 yields the most stable and consistent performance, temperature 0 performs worse overall, and temperature 0.5 shows higher variance.
+Many-shot prompting is more affected by temperature changes than zero-shot prompting.
 Overall performance
-Many-shot prompting led to stronger alignment with expert consensus than zero-shot prompting (see Figure 5).
-Accuracy rose from 0.44 to 0.57, and Cohen’s $\kappa$ nearly doubled (0.10 → 0.18), indicating more consistent agreement beyond chance.
-Macro-precision also improved slightly (0.40 → 0.42), suggesting fewer false positives, while macro-recall decreased (0.53 → 0.46), implying a more conservative stance that missed some expert-labelled cases.
-Despite this trade-off, the overall macro-F1 increased from 0.34 to 0.40, reflecting a net gain in balanced performance.
-Per-class performance: Strong and Moderate improve, Poor remains unreliable.
+Many-shot prompting achieved stronger alignment with expert consensus than zero-shot prompting (see Figure 6).
+Accuracy improved from 0.45 to 0.64, and Cohen’s $\kappa$ increased from 0.10 to 0.26.
+While Cohen’s $\kappa$ shows an improvement, it still remains within the slight-agreement range.
+Macro-precision also rose modestly (0.40 → 0.44), indicating fewer false positives, while macro-recall decreased slightly (0.54 → 0.50), suggesting a more conservative stance that misses some expert-labelled cases.
+Despite this trade-off, the overall macro-F1 increased from 0.35 to 0.44, representing a clear improvement in balanced performance.
+Per-class performance.
 At the category level, many-shot prompting yields improvements for Strong and Moderate cases but not for Poor.
-For Strong, recall rises (0.48→0.60) and F1 improves (0.41→0.49), indicating more reliable detection of clear joint attention.
-For Moderate, which dominates the dataset, both precision (0.80→0.83) and recall (0.42→0.57) increase, producing the largest F1 gain (0.55→0.67).
-By contrast, Poor segments remain highly challenging: recall drops (0.70→0.20) and F1 declines (0.07→0.04).
+For Strong, recall increases from 0.47 to 0.64 and F1 improves from 0.41 to 0.52, indicating more reliable detection of clear joint attention.
+For Moderate, which dominates the dataset, both precision (0.80→0.85) and recall (0.44→0.65) increase, producing the largest F1 gain (0.57→0.74).
+By contrast, Poor segments remain highly challenging: recall drops sharply (0.70→0.20) and F1 decreases (0.07→0.06).
 This reflects two issues.
-First, zero-shot prompting severely over-predicted poor cases, inflating recall but with almost no precision (0.04).
-many-shot prompting suppresses this over-prediction, yielding more conservative outputs but still very low precision (0.02).
-Second, the data distribution itself is highly imbalanced: across the dataset, Moderate accounts for 469 segments (73.5%), Strong for 136 (21.3%), while Poor is represented by only 10 examples (1.6%).
-Such scarcity makes it difficult to validate or learn reliable patterns for low-fidelity behaviours in a many-shot condition.
+First, zero-shot prompting substantially over-predicts Poor cases, producing high recall but extremely low precision (0.04).
+Many-shot prompting suppresses this over-prediction, yielding more conservative outputs but still very low precision (0.04→0.04).
+Second, the dataset distribution is highly imbalanced: Moderate accounts for 469 segments (73.5%), Strong for 136 (21.3%), and Poor for only 10 examples (1.6%).
+Such scarcity makes it difficult for the model to learn or validate reliable patterns for low-fidelity Poor behaviours, regardless of prompting method.
 Figure 6.
@@
 many-shot consistently improves accuracy, macro-F1, and Cohen’s $\kappa$ in the overall condition, while also boosting performance on Strong and Moderate categories.
-Performance on the Poor category remains weak under both conditions, reflecting its limited representation in the dataset.:
+Performance on the Poor category remains weak under both conditions, reflecting its limited representation in the dataset.
+The radar plot compares the performance of zero-shot (red) and many-shot (blue) prompting, showing that many-shot prompting consistently improves overall metrics and performance for the ’Strong’ and ’Moderate’ categories, while performance on the ’Poor’ category remains low due to its limited dataset representation.
 5. Discussion
 5.1. Greater Alignment at the Level of Behavioural Observations
-SLPs showed strong alignment with one another with behavioural observation, consistently pointing to gaze, action, and vocalisation as the central dimensions for describing joint attention.
-Our structured prompting method leveraged this shared consensus and achieved similarly high alignment: MLLMs could reliably capture and reproduce these behavioural cues.
-Because expert criteria were already aligned, observation proved to be a tractable entry point for human–AI alignment.
-By guiding MLLMs to “see” behaviours through these shared dimensions, we can achieve stable and reproducible outputs.
-This suggests that systems for analysing Parent-Child Interaction should prioritise clear, structured observation before attempting higher-level reasoning.
-The relative ease of alignment in observation highlights a broader design lesson: when experts already share common descriptors, MLLMs can be scaffolded to replicate them reliably.
-Multimodal models are particularly well-suited to capture low-level, directly perceivable cues—for instance, gaze direction, vocal onset, or gesture trajectories—because such behaviours are already described consistently across experts.
-Achieving alignment in these cases mainly requires iterative confirmation with human experts, rather than inventing new criteria.
-For example, in education (xing2024survey), teachers can consistently agree on whether a child raised a hand, spoke out loud, or looked at the teacher.
-In UX evaluation (vermeeren2010user), experts reliably agree on whether a user paused, clicked, or hovered over an interface element.
-In creative work critiques (alabood2023systematic), reviewers may all notice surface-level features such as colour saturation or the use of bold lines.
-These cues are easier for multimodal models to observe, making observation alignment relatively tractable.
+SLPs showed strong alignment with one another in behavioural observation, consistently pointing to gaze, action, and vocalisation as the central dimensions for describing joint attention.
+Our structured prompting method leveraged this shared consensus and achieved similarly high alignment.
+MLLMs could reliably capture and reproduce these behavioural cues.
+Because experts already converge on these descriptors, observation proved to be a tractable entry point for human–AI alignment.
+This tractability reflects a broader phenomenon.
+When the domain involves low-level, directly perceivable behaviours, experts tend to describe events in highly similar ways.
+In education (Xing et al., 2024), teachers reliably agree on whether a child raised a hand or spoke out loud.
+In UX evaluation (Vermeeren et al., 2010), experts consistently identify pauses, clicks, and hovers.
+In creative work critiques (Alabood et al., 2023), reviewers often converge on surface-level cues such as colour saturation or bold outlines.
+These shared descriptors create a stable grounding that multimodal models can be scaffolded to replicate.
+As a result, observation-level alignment can be achieved before attempting any higher-level judging.
+This suggests that systems for analysing Parent-Child Interaction should prioritise clear and structured observation as their first layer of alignment.
+By guiding MLLMs to “see” behaviours through common expert descriptors, we can produce stable and interpretable outputs that form a reliable substrate for subsequent judging tasks.
+Design Takeaway #1: Provide Perceptual Assistance With Transparent and Interpretable Cues.
+MLLMs can meaningfully support early-stage observation.
+In our study, the model achieved 81% accuracy in identifying low-level behavioural cues, showing strong alignment with SLPs in detecting gaze shifts, object-directed actions, and vocalisations.
+However, these machine-generated observations cannot replace human observation, especially in a clinically serious and nuanced behavioural assessment task.
+Even with relatively strong cue-level performance, SLPs should not simply trust model outputs.
+They must actively review, verify, and integrate these cues into their own observation workflow.
+These cues nevertheless remain valuable.
+They are precisely the behaviours that SLPs spend substantial time documenting during and after sessions, both for clinical records and to support later judgement.
+SLP2 specifically noted that gathering concrete behavioural examples is a time-consuming but essential part of subsequent assessment.
+By surfacing structured and timestamped observations, MLLMs can reduce documentation burden and provide a consistent baseline for longitudinal tracking, provided that SLPs remain the primary interpreters.
+This highlights a concrete design direction.
+Systems should help SLPs use model-extracted cues efficiently and reflectively.
+Interfaces should allow interactive editing of cue descriptions, timestamp correction, merging or splitting of events, and clinician-authored annotations.
+These features can transform model detections into editable professional artefacts.
+Transparent visualisations, such as layered timelines, collapsible cue categories, uncertainty indicators, and explicit evidence, can further support rapid verification while preventing over-reliance on automation.
+Through these mechanisms, perceptual assistance augments rather than replaces human expertise and embeds MLLM-derived cues into clinical judgement in a controlled and interpretable manner.
 5.2. Partial Alignment at the Level of Judgement
-Overall alignment scores at the judgement stage were modest, reflecting the difficulty of capturing interpretive nuances.
-Nevertheless, we found that many-shot prompting consistently outperformed zero-shot, suggesting that incorporating examples grounded in SLP expertise helps models approximate expert reasoning more closely.
-This indicates that expert knowledge does play a role in supporting judgement-level alignment, even if the gains are incremental.
-The results highlight the need for more deliberate strategies to improve judgement-level alignment.
-Few-shot prompting offers one pathway, but more sustained approaches such as fine-tuning with expert-labelled data or retrieval-augmented generation (RAG) may be necessary to embed richer interpretive cues.
-At the same time, because judgement involves a degree of subjectivity, future systems may need to actively explore how different forms of expertise shape interpretation, and provide mechanisms to surface or adapt to these perspectives rather than converge on a single fixed standard.
-This raises a broader question about how to handle inherently subjective layers of expertise in human-AI alignment.
-At the observational level, alignment is more attainable because experts tend to converge on concrete cues such as gaze, action, or vocalisation.
-At the judgement level, however, interpretations vary and are often context-dependent, making consensus elusive.
-This challenge is not unique to joint attention: expert decision-making in many fields is notoriously tacit and difficult to distil into explicit rules, drawing on years of situated practice and subtle interpretive skill.
-For instance, in medicine, radiologists may differ in judging whether a faint shadow on a scan indicates pathology or normal variation.
-In law, judges interpreting similar facts may reach different conclusions based on contextual reasoning.
-In Parent-Child Interaction analysis, where interpretive diversity is intrinsic, the goal may be less about maximising consensus and more about enabling systems to represent, contrast, and adapt to multiple plausible viewpoints.
-This perspective opens up space for HCI and AI research to treat subjectivity not as noise to be eliminated, but as a critical design consideration in the development of judgment-sensitive systems.
-In education, a raised hand may be judged by one teacher as evidence of engagement, while another views it as disruptive if out of turn.
-In UX evaluation, a pause may be seen by one expert as a usability breakdown, but by another as part of a normal learning curve.
-In creative work critiques, the same bold use of colour may be praised by one reviewer as innovative, yet criticised by another as unbalanced.
-These examples highlight that while observation alignment is attainable, judgement alignment may be less realistic or even undesirable.
-For exploratory domains like Parent-Child Interaction analysis, acknowledging and surfacing disagreement may be more valuable than enforcing consensus, as it allows users to reflect on alternative interpretations rather than rely on a single definitive output.
+Alignment at the judgement stage was modest, reflecting the difficulty of capturing interpretive nuances.
+Many-shot prompting consistently outperformed zero-shot prompting, suggesting that examples grounded in SLP expertise help the model approximate expert judging more closely.
+These gains were incremental, however, indicating that judgement-level alignment is shaped not only by prompting strategies but also by natural variation in expert perspectives.
+Judgement in Parent-Child Interaction analysis is inherently subjective and context dependent.
+Different SLPs weight behavioural cues differently, and their interpretations can diverge even when they observe the same interaction.
+Similar variation also appears in other expert domains.
+Teachers may interpret a raised hand as engagement, whereas others see it as disruption; UX evaluators may view a pause as confusion, while others treat it as normal exploration; and reviewers in creative critique often make contrasting assessments of the same work.
+These examples illustrate that judgement alignment is fundamentally different from observation alignment.
+In our dataset, heterogeneous annotations made it difficult to derive a single, coherent judgement space.
+We also observed pockets of higher agreement among certain experts, such as SLP1 and SLP2 ($\alpha=0.31$).
+When data are limited and expert perspectives vary widely, such internally consistent subgroups point to a possible direction for future work: examining whether alignment strategies that draw on individual experts or small groups could offer more stable early-stage models.
+Design Takeaway #2: Calibrate AI Judgement to Task Needs Through Expert-Aligned Adaptation Given the heterogeneous annotation landscape, a practical direction for HCI systems is to support personalised, expert-aligned adaptation.
+For SLP-facing tools, tailoring the AI to an individual clinician’s own judgement tendencies may be particularly useful, as their criteria for a given task tend to remain relatively stable over time.
+For parent-facing tools, however, a consensus-informed representation derived from multiple experts may provide clearer and more balanced guidance.
+Developing such a system responsibly would require substantially larger and more diverse datasets, including naturalistic parent–child interactions and broader SLP annotations, to ensure that any consensus model is reliable, generalisable, and appropriate for use in parent-oriented contexts.
+At the presentation level, SLPs also highlighted the importance of conveying joint attention as an overall pattern rather than a sequence of isolated segments.
+SLP3 suggested a colour-bar overview where each segment’s label is encoded as a colour across the entire timeline.
+Such visual summaries provide both SLPs and parents with an immediate, intuitive understanding of interaction quality and help them focus efficiently on segments that may require further attention.
+Across these use cases, AI outputs should be treated as supportive references rather than definitive judgements.
+From a reliability standpoint, the final assessment should still be made by an SLP.
+Judgement in Parent-Child Interaction is only one part of an SLP’s broader role; clinicians also provide therapy, coach parents, interact with children, and tailor guidance to highly specific family contexts.
+These relational and instructional responsibilities cannot be automated.
+AI systems can, however, enhance clinical efficiency by surfacing behavioural patterns, organising information, and providing structured summaries that clinicians can interpret and incorporate into therapeutic practice.
+Future work should investigate how personalised expert aligned models and consensus informed models might be integrated into broader clinical workflows and how these two modes could support different stakeholders within real developmental assessment and intervention pathways.
 5.3. Limitations
@@
 Notably, one SLP emphasised that detecting Poor joint attention is often clinically more critical, and suggested that recordings featuring neurodiverse children (e.g., autism-focused datasets) may not only be more representative of practice but also easier to annotate consistently.
-Expanding future datasets to include longer interactions, neurodiverse populations (benton2014diversity), and recordings from real-world home or school settings will be essential for improving ecological validity.
+Expanding future datasets to include longer interactions, neurodiverse populations (Benton et al., 2014), and recordings from real-world home or school settings will be essential for improving ecological validity.
+Human Editing
+A methodological limitation of our approach is that the input to the Stage 2 judge process was based on manually corrected descriptions rather than the raw outputs produced by the MLLM in Stage 1.
+Although fewer than 15% of all entries required correction and each modification followed a minimal-editing principle that preserved the subject–verb–object structure, these corrections removed factual errors that would otherwise propagate into Stage 2.
+Our intention was to isolate the effect of the judge process itself, but this design choice means that the Stage 2 results likely represent an upper bound on performance in a fully end-to-end setting.
+Future work will evaluate the judge process using uncorrected model outputs to obtain a more conservative estimate of system behaviour.
 Positioning and Scope
@@
 By comparing three SLP’ perspectives on joint attention with MLLM outputs, we identified a clear contrast: observation-level alignment was relatively robust, while judgement-level alignment remained elusive due to differences among experts themselves.
-These findings suggest that MLLMs can serve as reliable observers when scaffolded by shared behavioural cues, but that interpretive alignment requires more nuanced approaches and may benefit from exposing diverse perspectives rather than enforcing consensus.
+These findings suggest that MLLMs can serve as reliable observers when scaffolded by shared behavioural cues, and they highlight how future HCI systems may draw value from both individual or subgroup SLP judgements and areas of broader consensus across the expert cohort.
 Our contribution should be understood as a case-based investigation of alignment rather than a validated framework.